### v2

# Infinitely many genus two and three curves of small fixed positive rank

Abstract.
We prove that there exist infinitely many genus two curves over $\mathbb{Q}$ whose Jacobians have ranks equal to a fixed number between 00 and $11$, and genus three curves over $\mathbb{Q}$ whose Jacobians have ranks equal to a fixed number between 00 and $4$. Furthermore, we consider a generalisation of these results over number fields. We prove that for any number field $K$, there are infinitely many genus two curves over $K$ whose Jacobians have ranks over $K$ equal to each of 0, 1, or 2.
1. Introduction
Let $K$ be a number field, and $C/K$ a nice curve111smooth, projective, geometrically irreducible with its Jacobian $\mathrm{Jac}(C)$. Mordell-Weil theorem states that $\mathrm{Jac}(C)(K)$ is a finitely generated abelian group, i.e., isomorphic to $\mathbb{Z}^{r}\times G$, where $G$ is a finite group. The number $r$ is called the rank of $\mathrm{Jac}(C)$ over $K$.
Even though there are techniques to compute the rank $r$ in certain cases, for example, using descent [Sch95, Sto01, SS03], it is still a difficult problem to determine the exact value of $r$. In principle, the problem can be split into two tasks. The lower bound on $r$ can be obtained from finding enough $\mathbb{Z}$-linearly independent elements in $\mathrm{Jac}(C)(K)$. The upper bound on $r$ can be computed from using descent and computing Selmer groups of $\mathrm{Jac}(C)(K)$. When the two bounds are equal, then the exact value of $r$ can be determined.
The problem of computing the rank of $\mathrm{Jac}(C)(K)$ is of high significance. We give some of its important consequences, as provided below.
  • For studying the set $C(K)$, it is very useful to know the rank of $\mathrm{Jac}(C)(K)$, as in some cases a powerful method of Chabauty and Coleman [Cha41, Col85] can be applied to determine $C(K)$. Some generalisations of the method by Chabauty and Coleman, dependent on the rank of $\mathrm{Jac}(C)(K)$, can be found in [Kim05, Kim09, BDM+19, Sik09]. These methods have led to numerous results on studying modular curves that parametrise certain families of elliptic curves, their modularity, torsion subgroups, isogenies and residual Galois representations.
  • One part of the proof of the Hilbert 10th problem over the ring of integers $\mathcal{O}_{K}$, where $K$ is a number field, amounts to prove that for a given elliptic curve $E/\mathcal{O}_{K}$ of rank exactly 1, there exists its quadratic twist of rank also exactly 1, as proven recently for most number fields in [KP24]. One can also construct infinitely many elliptic curves of rank exactly 1 which are not $\overline{K}$-isomorphic, as proven independently in [Zyw25a].
  • The Birch-Swinnerton-Dyer conjecture [BSD65] proposes intricate relations between algebraic properties and analytic properties of abelian varieties. Finding evidence supporting the conjecture has harbored fruitful results in understanding arithmetic properties of elliptic curves [GZ86, Gro91, ZI04, DLR15].
In the case of modular curves, using Eichler-Shimura relations, it is possible to decompose their Jacobians and compute their ranks by computing the ranks of their components. Still, this technique is not possible for most of the curves. Even for hyperelliptic curves, for which we can try to apply the 2-descent, it is not clear if the upper bound obtained from descent will match the number of $\mathbb{Z}$-linearly independent elements in $\mathrm{Jac}(C)(K)$. Knowing the existence of additional $\mathbb{Z}$-linearly independent elements in $\mathrm{Jac}(C)(K)$ may not be helpful, because the 2-Selmer group of $\mathrm{Jac}(C)$ can be of large dimension, which is not sufficient to compute the rank of $\mathrm{Jac}(C)(K)$.
For example, in [Gaj22], the first named author presents infinite families of curves over $\mathbb{Q}$ whose Jacobians have large ranks. Namely if we denote by $g(C)$ the genus of $C$, then an infinite family of hyperelliptic curves such that $\mathrm{Rank}(J(\mathbb{Q}))\geq g(C)-1$ is given in [Gaj22, Theorem 12(1)], whereas $\mathrm{Rank}(J(\mathbb{Q}))\geq g(C)$ is given in [Gaj22, Theorem 12(2)]. The main idea was to find curves with too many rational points that would violate Coleman’s bound [Col85], or Stoll’s improvement [Sto06]. However, giving an exact rank of such curves seems extremely difficult.
In this manuscript, we shift our focus to studying arithmetic statistics of ranks of $\mathrm{Jac}(C)(K)$, where $C$ varies over an infinite family of curves. There is a wealth of previous studies which analyse the statistics of ranks over infinite family of curves, for example over universal families of elliptic curves [BS15], quadratic twist families of elliptic curves [KMR14, Smi22a, Smi22b, Smi25], and cubic twist families of elliptic curves [ABS22, KS24].
1.1. Our result and strategy
The goal of this manuscript is to construct infinitely many genus two and three curves of fixed positive rank.
In Section 2, we consider bielliptic genus two curves $C$ such that $\mathrm{Jac}(C)$ is isogenous to product of two elliptic curves $E_{1}$ and $E_{2}$. We divide the construction into 3 cases. In §2.1, we consider the case over number fields $K$ where both $E_{1}$ and $E_{2}$ do not have any $K$-rational $2$-torsion points. In §2.2, we examine elliptic curves over $\mathbb{Q}$ having marked rational $2$-torsion points. In §2.3, we use elliptic curves over $\mathbb{Q}$ having all rational $2$-torsion points. We then construct families of genus two curves $C$ by fixing one curve $E_{2}$ of the prescribed rank $r$, and quadratically twisting $E_{1}$ except in §2.4, which we explain later. This allows us to utilise results from previous literature (such as [Nag29, Fre84, KMR14]) that there are infinitely many such twists of rank 0 or 1 depending on the choice of $E_{1}$. We can hence construct infinitely many bielliptic curves of prescribed rank. For $K=\mathbb{Q}$, we construct the desired families for $0\leq r\leq 11$.
The strategy we take for genus three curves is similar. We consider bielliptic curves $C$ such that $\mathrm{Jac}(C)$ is isogenous to a product of an elliptic curve $E$ and the Jacobian of a genus two curve $C^{\prime}$. We then construct families of genus three curves $C$ by fixing the genus two curve $C$ of prescribed rank $r$, and quadratically twisting $E$. We then use previous studies [Kan13, Smi22a] to obtain that over $K=\mathbb{Q}$, there are infinitely many such twists of rank 0. This allows us to construct infinitely many bielliptic curves of ranks 00, $1$, and $2$ over $\mathbb{Q}$.
For most of our constructions, it is crucial to fix an elliptic curve of the prescribed rank and to aim for twists of rank 0. Other strategies, for instance, to fix an elliptic curve of rank 0 and to aim for infinitely many twists of rank $r$, seem much more difficult, as in principle, it is much easier to show the existence of infinite families of curves of rank 0 rather than positive rank. We prove that there are not many rational points on these curves, as explained in §4.1. We use positive rank twists in one of the examples in §2.2.
In §2.4, we also construct infinitely many curves over any number field $K$ of genus two, whose rank of the Jacobian over $K$ is each of 0, 1, and 2. These constructions use groundbreaking results in [Smi22b, KP25], which give an existence of infinitely many quadratic twists of a certain elliptic curve over $K$ of rank 0 [Smi22b] or 1 [KP25]. The conditions for elliptic curves for which their theorems can be applied are more complicated than those over $\mathbb{Q}$. Unlike the previous three subsections, the non-zero rank examples we construct crucially depend on rank 1 quadratic twist families. Checking whether the conditions appearing in [KP25] are applicable to our construction of genus two curves is an additional technicality we address in this subsection. The heart of our result for number fields lies in Theorem 2.8, where we propose an intricate construction for such genus two curves. Our work can be considered as a continuation of the recent progress by pioneering constructions of infinitely many elliptic curves of rank exactly equal to $1$ over any number field [KP25, Zyw25a]. More concretely, our work is an application of the results from [KP25] and of the same question for rank 0 [Smi22b].
We finally ask for an existence of infinitely many genus two or three curves defined over $\mathbb{Q}$ with a fixed positive rank of its Jacobian over $\mathbb{Q}$, but with an additional condition that the Jacobian is absolutely simple. This condition makes it a difficult question, see §4.2.
1.2. Explicit families of positive rank
Even though there are several instances of families of curves of fixed positive rank, such as [ABS22, KP25, Zyw25a], not so many give examples of families with explicit parametrisation. There are nevertheless some previous studies which give explicit parametrisation of families of elliptic curves over $\mathbb{Q}$ with rank $1$, such as those proven by Monsky [Mon92] and Dasgupta and Voight [DV18], those with rank $2$ as proven recently by Zywina [Zyw25b]. We give explicit examples of curves of genus two and a fixed rank between 0 and 4 in §2.2 and §2.3.
2. Genus two curves
We consider bielliptic curves $C$ of genus two; then $\mathrm{Jac}(C)\sim E_{1}\times E_{2}$. In our constructions we fix one curve, say $E_{1}$, and we vary $E_{2}$. Note that, in general, $E_{1}\times E_{2}$ is not isogenous to a Jacobian of a genus two curve, but in our cases, the curves $E_{1}$ and $E_{2}$ are “compatible” so it is. More formally, we can glue them so that one gets a bielliptic curve of genus 2, whose Jacobian is isogenous to $E_{1}\times E_{2}$. We provide the constructions for all possible rational two torsion subgroups when considered over $\mathbb{Q}$, i.e., $E_{1}[2](\mathbb{Q})\cong E_{2}[2](\mathbb{Q})\in\{\langle O\rangle,\mathbb{Z}/2\mathbb{Z},\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}\}$. However, our construction in §2.1 works over more general number fields.
2.1. No rational two-torsion
Theorem 2.1.
Let $K$ be a number field and $r\in\mathbb{Z}_{\geq 0}$. Assume that there is an elliptic curve $E\colon y^{2}=x^{3}+a^{2}$, where $a\in\mathcal{O}_{K}\setminus\mathcal{O}_{K}^{3}$ such that $\mathrm{Rank}(E(K))=r$ and $K(E[2])/K$ is an $S_{3}$ extension. There are infinitely many genus two curves $C_{a,m}\colon y^{2}=x^{6}+m^{3}a$ defined over $K$ such that $\mathrm{Rank}(\mathrm{Jac}(C)(K))=r$.
Proof.
We first construct hyperelliptic curves of genus two which are also bielliptic. Given any $D\in K^{\times}$, consider the three curves over $K$:
$\begin{split}C_{D}&\colon y^{2}=x^{6}+D,\\ E_{1,D}&\colon y^{2}=x^{3}+D,\\ E_{2,D}&\colon y^{2}=x^{3}+D^{2}.\end{split}$ (1)
Then there exists an isogeny $\phi\colon\mathrm{Jac}(C_{D})\to E_{1,D}\times E_{2,D}$. A direct application of [BD11, Theorem 3.2] shows that $C_{D}$ is isogenous to a product of two elliptic curves whose Weierstrass equations are $E^{\prime}_{1,D}\colon y^{2}=x^{3}+D$ and $E^{\prime}_{2,D}\colon y^{2}=x^{4}+Dx$. A change of variables given by $X:=\frac{D}{x}$ and $Y:=\frac{Dy}{x^{2}}$ gives the Weierstrass equation for $E_{2,D}$.
Using these curves, we construct our desired infinite family of hyperelliptic curves as follows. Fix an element $a\in\mathcal{O}_{K}\setminus\mathcal{O}_{K}^{3}$ satisfying the conditions of the statement of the theorem. Let $\mathcal{D}_{a}$ be the collection of elements in $\mathcal{O}_{K}$ defined as
[…]
Then for any $D\in\mathcal{D}_{a}$, the Weierstrass equation for $E_{2,D}$ can be rewritten as $E\colon y^{2}=x^{3}+a^{2}$, which is independent of the choice of $D$. Because $K(E[2])/K$ is an $S_{3}$ extension, $K(E_{1,D}[2])/K$ is also an $S_{3}$ extension. We note that $\{E_{1,D}\}_{D\in D_{a}}$ is a family of quadratic twists of the elliptic curve $E^{\prime}\colon y^{2}=x^{3}+a$. By [KMR14, Theorem A], there exists an infinite subset $\mathcal{D}^{0}_{a}\subset\mathcal{D}_{a}$ such that $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(E_{1,D}/K)=0$, hence $\mathrm{Rank}E_{1,D}(K)=0$, for every $D\in\mathcal{D}^{0}_{a}$.
We note that for any $D\in\mathcal{D}_{a}$ there are infinitely many $D^{\prime}\in\mathcal{D}_{a}$ such that $C_{D}\not\cong C_{D^{\prime}}$. Indeed, suppose one chooses $D^{\prime}$ such that $DD^{\prime}\not\in\mathcal{O}_{K}^{2}$ and the number of prime ideals of $K$ dividing $D^{\prime}$ that are coprime to $(6a)\subset\mathcal{O}_{K}$ is strictly greater than that of $D$. Then there exists a place of $K$ such that $C_{D^{\prime}}$ has bad reduction whereas $C_{D}$ does not. Hence, we have that $\{C_{D}\}_{D\in D^{0}_{a}}$ is an infinite set of hyperelliptic curves such that $\mathrm{Rank}(\mathrm{Jac}(C_{D})(K))=r$. ∎
Corollary 2.2.
Let $0\leq r\leq 11$ be a fixed integer. There are infinitely many genus two curves $C$ defined over $\mathbb{Q}$, such that $\mathrm{Rank}(\mathrm{Jac}(C)(\mathbb{Q}))=r$.
Proof.
For $0\leq r\leq 4$, we apply Theorem 2.1 for the following values of $a=2,3,15,427,13\cdot 19\cdot 23\cdot 43$ for $r=0,1,2,3,4$, respectively. The corresponding ranks can easily be checked in Magma [BCP97]. For $5\leq r\leq 11$, we apply the main result from [ER04]. We note that the elliptic curve $y^{2}=x^{3}-432k^{2}$ is $3$-isogenous to the elliptic curve $y^{2}=x^{3}+16k^{2}$ [CP09, Section 1]. We can then set $a=4k$ for the following values of $k$ to produce infinitely many genus two curves of rank $r$ as provided below.
$\begin{split}r=5&:k=3\cdot 7\cdot 11\cdot 13\cdot 163,\\ r=6&:k=3\cdot 73\cdot 103\cdot 439,\\ r=7&:k=3\cdot 13\cdot 19\cdot 41\cdot 139\cdot 271,\\ r=8&:k=2\cdot 3\cdot 5\cdot 7\cdot 11\cdot 13\cdot 17\cdot 29\cdot 41\cdot 47\cdot 59,\\ r=9&:k=2\cdot 5\cdot 37\cdot 41\cdot 53\cdot 73\cdot 1231\cdot 4831,\\ r=10&:k=2\cdot 3\cdot 5\cdot 7\cdot 23\cdot 31\cdot 37\cdot 43\cdot 83\cdot 109\cdot 151\cdot 421,\\ r=11&:k=3\cdot 5\cdot 7\cdot 13\cdot 19\cdot 23\cdot 31\cdot 43\cdot 59\cdot 61\cdot 73\cdot 79\cdot 103\cdot 109\cdot 157\cdot 457.\end{split}$ (3)
Remark 2.3.
The condition of $K(E[2])/K$ being an $S_{3}$ extension amounts to $\zeta_{3}\notin K$ (necessary condition) and $3]{a}\notin K$. Therefore, for all number fields $K$ such that $3\nmid[K:\mathbb{Q}]$ and $\zeta_{3}\notin K$, this condition is satisfied. Otherwise, for fields with $\zeta_{3}\notin K$, but with $3]{a}\in K$, one can try to find another elliptic curve $E^{\prime}\colon y^{2}=x^{3}+a^{\prime 2}$ satisfying the conditions of Theorem 2.1.
Remark 2.4.
Suppose $K=\mathbb{Q}$. Using the main results from [Smi22a, Smi22b, Smi25], we can nullify the condition from Theorem 2.1 that $K(E[2])/K$ is an $S_{3}$ extension. Let $a\in\mathbb{Q}^{\times}$ be a fixed number. As before, let $E:y^{2}=x^{3}+a^{2}$. Then at least $50\%$ of genus two curves in $\{C_{D}\}_{D\in\mathcal{D}_{a}}$ have Jacobians of rank equal to $\mathrm{Rank}(E(\mathbb{Q}))$, and $100\%$ of genus two curves in $\{C_{D}\}_{D\in\mathcal{D}_{a}}$ have Jacobians of rank at most $\mathrm{Rank}(E(\mathbb{Q}))+1$.
2.2. Partial rational two-torsion
Now we consider the family of bielliptic curves $C_{m,d}\colon y^{2}=d^{3}x^{6}+m^{3}$, where $d,m\in\mathbb{Z}$ are squarefree. Similarly as before, [BD11, Theorem 3.2] implies that $\mathrm{Jac}(C_{m,d})$ is isogenous to $E_{d}\times E_{m}$, where $E_{d}\colon y^{2}=x^{3}+d^{3}$ and $E_{m}\colon y^{2}=x^{3}+m^{3}$, which is a quadratic twist of $E_{d}$ by $\frac{m}{d}$. Now, we fix $d$ and vary $m$, obtaining the following theorem.
Theorem 2.5.
Let $r\in\mathbb{Z}_{\geq 0}$ and $d\in\mathbb{Z}$ be such that $\mathrm{Rank}(E_{d}(\mathbb{Q}))=r$, where $E_{d}\colon y^{2}=x^{3}+d^{3}$. Then for the following curves of genus two $C_{p,d}\colon y^{2}=d^{3}x^{6}+p^{3}$, where $p$ is a prime number, we have
  • if $p\equiv 5\pmod{12}$, $\mathrm{Rank}(C_{p,d}(\mathbb{Q}))=r$;
  • if $p\equiv 3\pmod{4}$ and $p>3$, $\mathrm{Rank}(C_{p,d}(\mathbb{Q}))=r+1$;
Proof.
The proof follows from the text above and the fact that $\mathrm{Rank}(E_{p}(\mathbb{Q}))=0,1$, when $p\equiv 5\pmod{12}$, $p\equiv 3\pmod{4}$ and $p>3$, respectively, see [Fre84, P. 73, Satz 2, 3]. ∎
We note that this approach works only over $\mathbb{Q}$, but unlike the previous approach, it gives explicit families of infinitely many non-isomorphic curves of fixed positive rank. This family is an infinite set because the curve $C_{p,d}$ has a place of bad reduction at $p$. A quick Magma search gives us the following rank information:
  • $d=1$: $E_{1}\colon y^{2}=x^{3}+1$, $\mathrm{Rank}(E_{1}(\mathbb{Q}))=0$;
  • $d=2$: $E_{2}\colon y^{2}=x^{3}+2^{3}$, $\mathrm{Rank}(E_{2}(\mathbb{Q}))=1$;
  • $d=37$: $E_{37}\colon y^{2}=x^{3}+37^{3}$, $\mathrm{Rank}(E_{37}(\mathbb{Q}))=2$;
  • $d=506$: $E_{506}\colon y^{2}=x^{3}+506^{3}$, $\mathrm{Rank}(E_{506}(\mathbb{Q}))=3$.
For example, $y^{2}=506^{3}x^{6}+p^{3}$, where $p\equiv 3\pmod{4}$ and $p>3$, is a family of curves whose Jacobian has rank 4, whereas if $p\equiv 5\pmod{12}$, then this is a family of curves whose Jacobian has rank 3.
2.3. Full rational two-torsion - over $\mathbb{Q}$
Now we consider elliptic curves $E_{1},E_{2}/\mathbb{Q}$ such that $E_{1}[2](\mathbb{Q})\cong E_{2}[2](\mathbb{Q})\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. Since we want to vary one of them, we will specify one of them to be a congruent problem elliptic curve, i.e., let $E_{1}$ belong to the family $E^{p}\colon y^{2}=x^{3}-p^{2}x$, where $p\equiv 3\pmod{8}$; then by [Nag29], we have $\mathrm{Rank}(E^{p}(\mathbb{Q}))=0$. We want to use [BD11, Theorem 3.2], so, we look at the other elliptic curve of the form ($k\neq-1,0,1$)
$E_{2}\colon y^{2}=d(x-kp)(x^{3}-p^{2}x),$
which after replacing $(x,y)$ by $(px,p^{2}y)$ becomes $E_{2,d,k}\colon y^{2}=d(x-k)(x^{3}-x)$. From now on, we can search directly in Magma for curves $E_{2,d,k}$ of any rank we want. Still, if we prefer to use the standard way of expressing elliptic curves, a linear change of variables leads to the following model
$E^{\prime}_{2,d,k}\colon dy^{2}=(x+k^{2}-k)(x+k^{2}-1)(x+k^{2}+k).$
We obtain the following theorem, which immediately follows from [BD11, Theorem 3.2].
Theorem 2.6.
Let $r\in\mathbb{Z}_{\geq 0}$, $k\in\mathbb{Z}\setminus\{-1,0,1\}$, and squarefree $d\in\mathbb{Z}$ be such that $\mathrm{Rank}(E^{\prime}_{2,d,k}(\mathbb{Q}))=r$, where $E^{\prime}_{2,d,k}\colon dy^{2}=(x+k^{2}-k)(x+k^{2}-1)(x+k^{2}+k)$. Then, for the following curves of genus two
$C_{d,k,p}\colon dy^{2}=x^{6}+3dkpx^{4}+(3k^{2}p^{2}-1)d^{2}x^{2}+(k^{3}p^{3}-kp)d^{3},$
where $p\equiv 3\pmod{8}$ is a prime number, we have $\mathrm{Rank}(C_{d,k,p}(\mathbb{Q}))=r$.
We note that given a fixed value of $k$ and $d$, the curve $C_{d,k,p}$ has a place of bad reduction at $p$. Hence, the collection of curves $\{C_{d,k,p}\}_{p\equiv 3\mod 8}$ is an infinite set. Here also we obtain results over $\mathbb{Q}$, but we can give some interesting explicit examples. Testing the curves $E^{\prime}_{2,d,k}$ for $d=1$ in Magma, we obtain the following rank information:
  • $k=2$: $\mathrm{Rank}(E^{\prime}_{2,1,2}(\mathbb{Q}))=0$;
  • $k=3$: $\mathrm{Rank}(E^{\prime}_{2,1,3}(\mathbb{Q}))=1$;
  • $k=11$: $\mathrm{Rank}(E^{\prime}_{2,1,11}(\mathbb{Q}))=2$;
  • $k=43$: $\mathrm{Rank}(E^{\prime}_{2,1,43}(\mathbb{Q}))=3$;
  • $k=329$: $\mathrm{Rank}(E^{\prime}_{2,1,329}(\mathbb{Q}))=4$.
For example, one more explicit family of curves of genus two whose Jacobians have rank 4 is the family $C_{1,329,p}$, where $p\equiv 3\pmod{8}$.
2.4. Full rational two-torsion - over any number field
Let $K$ be any number field. We now consider elliptic curves $E_{1},E_{2}/K$ such that $E_{1}[2](K)\cong E_{2}[2](K)\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. Unlike the case where $K=\mathbb{Q}$, we will vary both elliptic curves over their families of quadratic twists. Let $a_{1},a_{2},a_{3}\in\mathcal{O}_{K}\setminus\{0\}$. We consider the following elliptic curve, where $m$ is any square-free element of $\mathcal{O}_{K}$.
$E_{1,m}$ $:y^{2}=m(x-a_{1})(x-a_{2})(x-a_{3})$
We then choose $\alpha\in\mathcal{O}_{K}\setminus\{a_{1},a_{2},a_{3}\}$ and a square-free element $d\in\mathcal{O}_{K}$ to define the following elliptic curve by using [BD11, Theorem 3.2]:
$E_{2}:y^{2}=d(x-m\alpha)(x-ma_{1})(x-ma_{2})(x-ma_{3}).$
After replacing $(x,y)$ with $(mx,m^{2}y)$, we obtain $E_{2,d}:y^{2}=d(x-\alpha)(x-a_{1})(x-a_{2})(x-a_{3})$. We can summarize our findings as stated in the following theorem.
Theorem 2.7.
Let $r_{1},r_{2}\in\mathbb{Z}_{\geq 0}$, $a_{1},a_{2},a_{3}\in\mathcal{O}_{K}$, $\alpha\in\mathcal{O}_{K}\setminus\{a_{1},a_{2},a_{3}\}$, and $m,d\in\mathcal{O}_{K}$ be square-free. Let $E_{1,m}:y^{2}=m(x-a_{1})(x-a_{2})(x-a_{3})$ and $E_{2,d}:y^{2}=d(x-\alpha)(x-a_{1})(x-a_{2})(x-a_{3})$ be elliptic curves such that $\mathrm{Rank}(E_{1,m}(K))=r_{1}$ and $\mathrm{Rank}(E_{2,d}(K))=r_{2}$. Then for the following curve of genus two
$C_{m,d}:y^{2}=m\left(\frac{1}{d}x^{2}+\alpha-a_{1}\right)\left(\frac{1}{d}x^{2}+\alpha-a_{2}\right)\left(\frac{1}{d}x^{2}+\alpha-a_{3}\right)$
we have $\mathrm{Rank}(C_{m,d}(K))=r_{1}+r_{2}$.
Combining the above theorem with recent breakthroughs in distribution of algebraic ranks of quadratic twist families of elliptic curves over number fields [Smi22b, KP25], we can prove the following result for genus 2 curves over any number field $K$.
Theorem 2.8.
Let $K$ be any number field. Let $0\leq r\leq 2$ be a fixed integer. Then there are infinitely many genus two curves $C$ defined over $K$ such that $\mathrm{Rank}(Jac(C)(K))=r$.
Proof.
We choose nine prime elements $p_{i},q_{i},r_{i}$ of $\mathcal{O}_{K}$ so that the following conditions are satisfied:
  • $p_{i},q_{j},r_{k}$ are totally positive.
  • $p_{i},q_{j},r_{k}\equiv 1\mod 8\mathcal{O}_{K}$.
  • #$\#\mathcal{O}_{K}/(p_{i}),\#\mathcal{O}_{K}/(q_{i}),\#\mathcal{O}_{K}/(r_{i})>5$.
  • For every $1\leq i,j,k\leq 3$, we have $\left(\dfrac{p_{i}}{q_{j}}\right)=\left(\dfrac{q_{j}}{r_{k}}\right)=\left(\dfrac{r_{k}}{p_{i}}\right)=1$.
We will use the following result which is an immediate consequence of [Mit56, P. 20, Theorem]
Lemma 2.9.
Let $I$ be an ideal of $K$ and $\rho\in\mathcal{O}_{K}$ be totally positive such that the ideals $I$ and $(\rho)$ are coprime. Then there are infinitely many totally positive prime elements of $p\in K$ satisfying $p\equiv\rho\pmod{I}$.
Now we apply Lemma 2.9 to the ideal $I=(8)$ and $\rho=1$ to find totally positive prime elements $p_{1},p_{2},p_{3}\in\mathcal{O}_{K}$ such that #$\#\mathcal{O}_{K}/(p_{i})>5$. Then, we apply Lemma 2.9 to the ideal $I=(8p_{1}p_{2}p_{3})$ and $\rho=1$ to find totally positive prime elements $q_{1},q_{2},q_{3}\in\mathcal{O}_{K}$ such that #$\#\mathcal{O}_{K}/(q_{i})>5$. Finally, we apply Lemma 2.9 to the ideal $I=(8p_{1}p_{2}p_{3}q_{1}q_{2}q_{3})$ and $\rho=1$ to find totally positive prime elements $r_{1},r_{2},r_{3}\in\mathcal{O}_{K}$ such that #$\#\mathcal{O}_{K}/(r_{i})>5$. By construction, we have that $\left(\dfrac{p_{i}}{q_{j}}\right)=\left(\dfrac{q_{j}}{r_{k}}\right)=\left(\dfrac{r_{k}}{p_{i}}\right)=1$.
For such a choice of $p_{i},q_{j},r_{k}$, we define $P$, $Q$, $R$ and $a_{1},a_{2},a_{3}$ as follows:
$P:=p_{1}p_{2}p_{3},\hskip 5.0ptQ:=q_{1}q_{2}q_{3},\hskip 5.0ptR:=r_{1}r_{2}r_{3},\hskip 5.0pta_{1}:=P^{2}Q,\hskip 5.0pta_{2}:=Q^{2}R,\hskip 5.0pta_{3}:=R^{2}P.$
Let $m,d$ be square-free elements of $\mathcal{O}_{K}$. We consider the family of genus $2$ curves $C_{m,d}$ defined as
$C_{m,d}:y^{2}=m\left(\frac{1}{d}x^{2}-a_{1}\right)\left(\frac{1}{d}x^{2}-a_{2}\right)\left(\frac{1}{d}x^{2}-a_{3}\right),$
which is isogenous to a product of two elliptic curves provided below:
$E_{1,m}$ $:y^{2}=m(x-a_{1})(x-a_{2})(x-a_{3}),$
$E_{2,d}$ $:y^{2}=d(x-a_{2}a_{3})(x-a_{1}a_{3})(x-a_{1}a_{2}).$
The Weierstrass equation for $E_{2,d}$ is obtained as follows. We set $\alpha=0$ and apply Theorem 2.7 to get $E_{2}:y^{2}=dx(x-a_{1})(x-a_{2})(x-a_{3})$. We then apply the rational change of coordinates which replaces $(x,y)$ with $(-\frac{1}{x}\cdot{a_{1}a_{2}a_{3}},\frac{y}{x^{2}}\cdot a_{1}a_{2}a_{3})$ to get the Weierstrass equation for $E_{2,d}$.
Notice that the four conditions on nine prime elements $p_{i},q_{j},r_{k}$ imply the following (because they all have odd valuation with respect to at least one of the primes $p_{i},q_{j},r_{k}$).
  • None of $(a_{1}-a_{2})(a_{1}-a_{3})$, $(a_{2}-a_{1})(a_{2}-a_{3})$, and $(a_{3}-a_{1})(a_{3}-a_{2})$ are squares in $K$.
  • None of $a_{2}a_{3}(a_{2}-a_{1})(a_{3}-a_{1})$, $a_{1}a_{3}(a_{1}-a_{2})(a_{3}-a_{2})$, and $a_{1}a_{2}(a_{1}-a_{3})(a_{2}-a_{3})$ are squares in $K$.
By [Smi22b, Theorem 2.14, Example 3.1], there exists infinitely many (in fact a positive proportion of) square-free elements $m,d\in\mathcal{O}_{K}$ such that $\mathrm{Rank}(E_{1,m}(K))=\mathrm{Rank}(E_{2,d}(K))=0$. For such choices of $m,d\in\mathcal{O}_{K}$ we have $\mathrm{Rank}(\mathrm{Jac}(C_{m,d})(K))=0$.
We turn our focus to constructing infinitely many curves $C_{m,d}$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{m,d})(K))=1$. By our choice of nine prime elements $p_{i},q_{j},r_{k}$ the elliptic curve $E_{1,1}$ satisfies the following three conditions:
  • $a_{1}-a_{2}$ has odd valuation with respect to primes $q_{1},q_{2}$, and $q_{3}$, whereas $a_{1}-a_{3}$ and $a_{2}-a_{3}$ are invertible squares modulo $q_{1},q_{2}$, or $q_{3}$.
  • $a_{2}-a_{3}$ has odd valuation with respect to primes $r_{1},r_{2}$, and $r_{3}$, whereas $a_{2}-a_{1}$ and $a_{3}-a_{1}$ are invertible squares modulo $r_{1},r_{2}$, or $r_{3}$.
  • $a_{3}-a_{1}$ has odd valuation with respect to primes $p_{1},p_{2}$, and $p_{3}$, whereas $a_{3}-a_{2}$ and $a_{1}-a_{2}$ are invertible squares $p_{1},p_{2}$, or $p_{3}$.
These three conditions are in fact the “3-generic” conditions appearing in [KP25, Definition 1.2]. By [KP25, Theorem 1.3], there exist infinitely many square-free elements $m$ such that $\mathrm{Rank}(E_{1,m}(K))=1$. We can again use [Smi22b, Theorem 2.14, Example 3.1] to obtain that there are infinitely many square-free elements $d\in\mathcal{O}_{K}$ such that $\mathrm{Rank}(E_{2,d}(K))=0$. For such $m$ and $d$ we have $\mathrm{Rank}(\mathrm{Jac}(C_{m,d})(K))=1$.
Lastly, we construct infinitely many curves $C_{m,d}$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{m,d})(K))=2$. We still choose nine totally positive prime elements $p_{i},q_{j},r_{k}\in\mathcal{O}_{K}$ in the same manner.
We also choose the same $a_{1},a_{2},a_{3}$, but now we change $\alpha$. Consider $\alpha\in\mathcal{O}_{K}$ such that $\alpha$ is a quadratic residue modulo all $p_{i},q_{j},r_{k}$, but such that $\alpha\not\equiv a_{1}\pmod{r_{i}}$, $\alpha\not\equiv a_{2}\pmod{p_{i}}$, and $\alpha\not\equiv a_{3}\pmod{q_{i}}$ for $1\leq i\leq 3$, by the assumption that all residue fields have cardinality greater than 5, we can certainly satisfy the conditions modulo each prime, and then by the Chinese Remainder Theorem obtain $\alpha$.
For such a choice of $p_{i},q_{j},r_{k},$ and $\alpha$, we consider the family of genus 2 curves $C_{m,d}$ defined as
$C_{m,d}:y^{2}=m\left(\frac{1}{d}x^{2}+\alpha-a_{1}\right)\left(\frac{1}{d}x^{2}+\alpha-a_{2}\right)\left(\frac{1}{d}x^{2}+\alpha-a_{2}\right)$
which is isogenous to a product of two elliptic curves provided below:
$E_{1,m}$ $:y^{2}=m(x-a_{1})(x-a_{2})(x-a_{3}),$
$E_{2,d}$ $:y^{2}=d(x+(a_{2}-\alpha)(a_{3}-\alpha))(x+(a_{1}-\alpha)(a_{3}-\alpha))(x+(a_{1}-\alpha)(a_{2}-\alpha)).$
As before, our choices of $p_{i},q_{j},r_{k}$ imply that $E_{1,1}$ is $3$-generic. The choice of $\alpha$ implies that the following three conditions hold:
  • $(\alpha-a_{3})(a_{1}-a_{2})$ has odd valuation with respect to primes $q_{1},q_{2}$, and $q_{3}$, whereas $(\alpha-a_{2})(a_{1}-a_{3})$ and $(\alpha-a_{1})(a_{2}-a_{3})$ are invertible squares modulo $q_{1},q_{2}$, or $q_{3}$.
  • $(\alpha-a_{1})(a_{2}-a_{3})$ has odd valuation with respect to primes $r_{1},r_{2}$, and $r_{3}$, whereas $(\alpha-a_{3})(a_{2}-a_{1})$ and $(\alpha-a_{2})(a_{3}-a_{1})$ are invertible squares modulo $r_{1},r_{2}$, or $r_{3}$.
  • $(\alpha-a_{2})(a_{3}-a_{1})$ has odd valuation with respect to primes $p_{1},p_{2}$, and $p_{3}$, whereas $(\alpha-a_{1})(a_{3}-a_{2})$ and $(\alpha-a_{3})(a_{1}-a_{2})$ are invertible squares modulo $p_{1},p_{2}$, or $p_{3}$.
These three conditions show that $E_{2,1}$ is 3-generic. By [KP25, Theorem 1.3], there exist infinitely many square-free elements $m$ and $d$ such that $\mathrm{Rank}(E_{1,m}(K))=1$ and $\mathrm{Rank}(E_{1,d}(K))=1$. For such $m$ and $d$ we have $\mathrm{Rank}(\mathrm{Jac}(C_{m,d})(K))=2$. ∎
3. Genus three curves
Similar to the genus two case, we consider bielliptic curve $C_{3}$ of genus $3$ Its Jacobian variety decomposes as $\mathrm{Jac}(C_{3})\sim E\times\mathrm{Jac}(C_{2})$, where $C_{2}$ is a genus two curve over $K$. Instead of working over general number fields, we restrict our focus to $K=\mathbb{Q}$. In a similar spirit to Theorem 2.1, we show that as long as there exists a curve $C_{2}$ with full rational 2-torsion whose Jacobian has algebraic rank $r$ one can vary $E$ to construct infinitely many curves of genus three and rank $r$.
Theorem 3.1.
Let $r\in\mathbb{Z}_{\geq 0}$. Let $a,b,c,d\in\mathbb{Q}^{\times}$ be distinct rational numbers. Assume that there is a hyperelliptic curve $C_{2}:y^{2}=x(x-a)(x-b)(x-c)(x-d)$ such that $\mathrm{Rank}(\mathrm{Jac}(C)(\mathbb{Q}))=r$. Then there are infinitely many genus three curves $C_{3,D}:Dy^{2}=(x^{2}-Da)(x^{2}-Db)(x^{2}-Dc)(x^{2}-Dd)$ defined over $\mathbb{Q}$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{3,D})(\mathbb{Q}))=r$.
Proof.
Given any $a^{\prime},b^{\prime},c^{\prime},d^{\prime}\in\mathbb{Q}^{\times}$ and a square-free integer $D\in\mathbb{Z}$, consider the following three curves over $\mathbb{Q}$.
$\begin{split}C_{3,D}&:Dy^{2}=(x^{2}-a^{\prime})(x^{2}-b^{\prime})(x^{2}-c^{\prime})(x^{2}-d^{\prime})\\ E_{D}&:Dy^{2}=(x+(d^{\prime}-a^{\prime})(d^{\prime}-b^{\prime}))(x+(d^{\prime}-a^{\prime})(d^{\prime}-c^{\prime}))(x+(d^{\prime}-b^{\prime})(d^{\prime}-c^{\prime}))\\ C_{2,D}&:Dy^{2}=x(x-a^{\prime})(x-b^{\prime})(x-c^{\prime})(x-d^{\prime}).\end{split}$ (4)
Then there exists an isogeny $\phi:\mathrm{Jac}(C_{3,D})\to E_{D}\times\mathrm{Jac}(C_{2,D})$, which follows directly from [Sha15, Theorem 1]. We note that the theorem is proven for genus three curves over $\mathbb{C}$, which can be generalized to genus three curves over $\mathbb{Q}$ we consider in this proof, see the end of [Sha15, Introduction]. The change of variables one needs to choose can be given as follows. For the elliptic curve $E_{D}$, we use the change of variables $\tilde{x}=x^{2}$ and $\tilde{y}=y$ to obtain Weierstrass equation for $E^{\prime}_{D}:D\tilde{y}^{2}=(\tilde{x}-a^{\prime})(\tilde{x}-b^{\prime})(\tilde{x}-c^{\prime})(\tilde{x}-d^{\prime})$. Afterwards, we use $X:=\frac{1}{(\tilde{x}-d^{\prime})(d^{\prime}-a^{\prime})(d^{\prime}-b^{\prime})(d^{\prime}-c^{\prime})}$ and $Y:=\tilde{y}\tilde{x}^{2}(d^{\prime}-a^{\prime})^{2}(d^{\prime}-b^{\prime})^{2}(d^{\prime}-c^{\prime})^{2}$ to obtain the Weierstrass equation for $E_{D}$. For the hyperelliptic curve $C_{2,D}$ we use the change of variables $X:=x^{2}$ and $Y:=xy$. We hence obtain that the function field of $C_{3,D}$, which is $\mathbb{Q}(x,y)$, contains a subfield $\mathbb{Q}(x^{2},y)$ corresponding to $E_{D}$ and another subfield $\mathbb{Q}(x^{2},xy)$ corresponding to $C_{2,D}$.
We now construct our desired infinite family of genus 3 curves. Let $a,b,c,d$ be integers which satisfy the following equation:
[…]
With this choice of parameters, the Weierstrass equations for the three curves can be rewritten as:
$\begin{split}C_{3,D}&:Dy^{2}=(x^{2}-Da)(x^{2}-Db)(x^{2}-Dc)(x^{2}-Dd),\\ E_{D}&:Dy^{2}=(x+(d-a)(d-b))(x+(d-a)(d-c))(x+(d-b)(d-c)),\\ C_{2}&:y^{2}=x(x-a)(x-b)(x-c)(x-d).\end{split}$ (6)
We note that the collection $\{C_{3,D}\}_{D\text{ square-free}}$ contains an infinite number of non-isomorphic genus $3$ curves over $\mathbb{Q}$, because as long as $D$ is coprime to $(d-a)(d-b)(d-c)(a-b)(b-c)(c-a)$, the elliptic curve $E_{D}$ has additive reduction over every prime divisors of $D$. By [Smi25, Theorem 1.1], there are infinitely many (in fact half of the) square-free integers $D$ such that $E_{D}$ satisfies $\mathrm{Rank}(E(\mathbb{Q}))=0$. (In fact, if one can guarantee that none of the triples of integers $(d-a)(d-b)(c-b)(c-a)$, $(d-a)(d-c)(b-c)(b-a)$, and $(d-b)(d-c)(a-c)(a-b)$ are squares, then one can use [Kan13] to show that a positive proportion of squarefree integers $D$ satisfies $\mathrm{Rank}(E_{D}(\mathbb{Q}))=0$). This implies that for infinitely many such choices of $D$, we have $\mathrm{Rank}(\mathrm{Jac}(C_{3,D})(\mathbb{Q}))=r$. ∎
As a consequence, we prove that there are infinitely many genus three hyperelliptic curves defined over $\mathbb{Q}$, whose rank of the Jacobian over $\mathbb{Q}$ is a fixed number between 0 and 4. We are especially thankful to Jennifer Balakrishnan for providing us with examples of genus two curves with full rational two torsion, whose Jacobians have algebraic rank equal to 3 and 4.
Corollary 3.2.
Let $0\leq r\leq 4$ be a fixed integer. Then there are infinitely many genus three curves $C$ defined over $\mathbb{Q}$ such that $\mathrm{Rank}(\mathrm{Jac}(C)(\mathbb{Q}))=r$.
Proof.
We apply Theorem 3.1 for the following values of $a,b,c,d$. Here we list examples for $0\leq r\leq 4$ in the table provided below. All the algebraic ranks can be checked using Magma [BCP97].
$r$ $a$ $b$ $c$ $d$
0 1 2 3 8
1 1 2 3 9
2 1 2 3 36
3 -50 -5 42 50
4 -50 1 27 50
4. Rational points on such curves and a challenge
4.1. Rational points on these curves
In this section, we only consider curves defined over $\mathbb{Q}$, so we exclude those from §2.4. Even though we construct curves of higher rank, it is easy to determine their rational points in most of our examples (except for $C_{p,d}$ from §2.2 when $p\equiv 3\pmod{4}$) because these curves admit a quotient map to a rank 0 elliptic curve. This elliptic curve has only finitely many points, and we can determine which pullbacks of these points are rational points. Now we focus on all but finitely many curves, which could still have more rational points.
For example, curves from §2.1 and §2.2 admit a quotient map to an elliptic curve of the form $E\colon y^{2}=x^{3}+s$, $s\in\mathbb{Z}$. By a well-known result [Sil09, Exercise 10.19], when $s\neq-432$ is not a square or a cube in $\mathbb{Z}$, then $E_{\mathrm{tors}}=\langle O\rangle$. This is the case for almost all curves in §2.1, so $C_{a,m}(\mathbb{Q})=\{\infty_{\pm}\}$ (with projective coordinates $(1:\pm 1:0)$).
On the other hand, also by [Sil09, Exercise 10.19], for $E_{p}\colon y^{2}=x^{3}+p^{3}$ with $p\equiv 5\pmod{12}$, we have $E_{p}(\mathbb{Q})=\{O,(-p,0)\}$. Hence, if $d$ in a square (i.e. $d=1$ since we assume $d$ is squarefree), then $C_{p,d}(\mathbb{Q})=\{\infty_{\pm}\}$. Otherwise, the set $C_{p,d}(\mathbb{Q})$ is empty. When $p\equiv 3\pmod{4}$, the problem of determining $C_{p,d}(\mathbb{Q})$ could be very difficult.
In §2.3, the structure of the $\mathbb{Q}$-rational points of rank 0 elliptic curve $E^{p}\colon y^{2}=x^{3}-p^{2}x$ is well known: $E^{p}(\mathbb{Q})=E^{p}(\mathbb{Q})[2]\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. Pullback of nontrivial two-torsion points to $C_{d,k,p}$ would give an $x$-coordinate satisfying $x^{2}\in\{-dkp,-d(k-1)p,-d(k+1)p\}$, but then $x\notin\mathbb{Q}$ for almost all $p$. Hence, if $d=1$, then we have $C_{d,k,p}(\mathbb{Q})=\{\infty_{\pm}\}$ almost always. Otherwise, we have $C_{d,k,p}(\mathbb{Q})=\emptyset$.
In Section 3, the rank 0 family of twists of elliptic curves is $E_{D}$ (using the notation from Section 3), for which we know $E_{D}(\mathbb{Q})[2]\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ and that it is a change of variables of $E^{\prime}_{D}$. Since twisting will not preserve rational coordinates of an $n$-torsion point, unless $n=2$, for almost all $D$, we will have that $E_{D}(\mathbb{Q})=E_{D}(\mathbb{Q})[2]$ (again, we recall that we consider such curves of rank 0). Pullback of these points correspond to the roots of the defining polynomial in $x$ of $E^{\prime}_{D}$ as well as those of $C_{3,D}$. Since we vary $D$, for almost all $D$ none of the $aD$, $bD$, $cD$, and $dD$ will be a square for fixed $a,b,c,d\in\mathbb{Z}$. We have $C_{3,D}(\mathbb{Q})=\emptyset$ almost always.
4.2. Challenge
If the Jacobian of a curve is decomposable as a product of Jacobians of curves, then we reduce the problem of computing the rank to smaller genus curves, which is, in principle, way easier. We believe that the problem becomes much harder if we restrict the question for curves $C$ of genus two such that $\mathrm{Jac}(C)$ is absolutely simple, which gives the final question.
Question 4.1.
Is there an infinite family of curves $C/\mathbb{Q}$ of genus two or three with absolutely simple Jacobian such that the rank of $\mathrm{Jac}(C)(\mathbb{Q})$ is equal to $1$ (or some other fixed positive integer)?
Acknowledgements
We would like to thank the Max Planck Institute for Mathematics for its generous support. We would like to thank Jennifer Balakrishnan, Dušan Dragutinović, Peter Koymans, Oana Padurariu, Carlo Pagano, Lazar Radičević, Alex Smith, and Wadim Zudilin for helpful discussions.