### v3

# Infinitely many hyperelliptic curves of small genus and small fixed rank, and of any genus and rank two

Abstract.
We prove that for any number field $K$ and any fixed genus $g\geq 2$, there are infinitely many non-isomorphic hyperelliptic curves of genus $g$ over $K$ whose Jacobians have rank over $K$ equal to each of 0, 1, or 2. As an example of our method, over $\mathbb{Q}$, we prove that there exist infinitely many non-isomorphic hyperelliptic curves of genus two, whose Jacobians have rank equal to a fixed number between $1$ and $11$, genus three and four curves with rank between $1$ and $4$, and genus five and six with rank between $1$ and $3$.
1. Introduction
Let $K$ be a number field, and $C/K$ a nice curve111smooth, projective, geometrically irreducible with its Jacobian $\mathrm{Jac}(C)$. Mordell-Weil theorem states that $\mathrm{Jac}(C)(K)$ is a finitely generated abelian group, i.e., isomorphic to $\mathbb{Z}^{r}\times G$, where $G$ is a finite group. The number $r$ is called the rank of $\mathrm{Jac}(C)$ over $K$.
Even though there are techniques to compute the rank $r$ in certain cases, for example, using descent [Sch95, Sto01, SS03], it is still a difficult problem to determine the exact value of $r$. In principle, the problem can be split into two tasks. The lower bound on $r$ can be obtained from finding enough $\mathbb{Z}$-linearly independent elements in $\mathrm{Jac}(C)(K)$. The upper bound on $r$ can be computed from using descent and computing Selmer groups of $\mathrm{Jac}(C)(K)$. When the two bounds are equal, then the exact value of $r$ can be determined.
The problem of computing the rank of $\mathrm{Jac}(C)(K)$ is of high significance. We give some of its important consequences, as provided below.
  • For studying the set $C(K)$, it is very useful to know the rank of $\mathrm{Jac}(C)(K)$, as in some cases a powerful method of Chabauty and Coleman [Cha41, Col85] can be applied to determine $C(K)$. Some generalisations of the method by Chabauty and Coleman, dependent on the rank of $\mathrm{Jac}(C)(K)$, can be found in [Kim05, Kim09, BDM+19, Sik09]. These methods have led to numerous results on studying modular curves that parametrise certain families of elliptic curves, their modularity, torsion subgroups, isogenies and residual Galois representations.
  • One part of the proof of the Hilbert 10th problem over the ring of integers $\mathcal{O}_{K}$, where $K$ is a number field, amounts to prove that there is an abelian variety $A/K$ and a quadratic extension $L/K$ such that $\mathrm{Rank}(A(L))=\mathrm{Rank}(A(K))>0$. This was first proven very recently in the groundbreaking work by Peter Koymans and Carlo Pagano [KP24] who constructed an elliptic curve $E/\mathcal{O}_{K}$ and a quadratic extension $L/K$ such that $\mathrm{Rank}(E(L))=\mathrm{Rank}(E(K))=1$. One can also construct infinitely many elliptic curves of rank exactly 1 which are not $\overline{K}$-isomorphic, as proven independently in [Zyw25b]. A different proof of Hilbert 10th problem using positive rank of Jacobians of some twist families of hyperelliptic curves was shown by another groundbreaking work by Levent Alpöge, Manjul Bhargava, Wei Ho, and Ari Shnidman [ABHS25].
  • The Birch-Swinnerton-Dyer conjecture [BSD65] proposes intricate relations between algebraic properties and analytic properties of abelian varieties. Finding evidence supporting the conjecture has harbored fruitful results in understanding arithmetic properties of elliptic curves [GZ86, Gro91, ZI04, DLR15].
In the case of modular curves, using Eichler-Shimura relations, it is possible to decompose their Jacobians and compute their rank by computing the rank of their components. Still, this technique is not possible for most of the curves. Even for hyperelliptic curves, for which we can try to apply the 2-descent, it is not clear if the upper bound obtained from descent will match the number of $\mathbb{Z}$-linearly independent elements in $\mathrm{Jac}(C)(K)$. Knowing the existence of additional $\mathbb{Z}$-linearly independent elements in $\mathrm{Jac}(C)(K)$ may not be helpful, because the 2-Selmer group of $\mathrm{Jac}(C)$ can be of large dimension, which is not sufficient to compute the rank of $\mathrm{Jac}(C)(K)$.
For example, in [Gaj22], the first named author presents infinite families of curves over $\mathbb{Q}$ whose Jacobians have large rank. Namely if we denote by $g(C)$ the genus of $C$, then an infinite family of hyperelliptic curves such that $\mathrm{Rank}(\mathrm{Jac}(C)(\mathbb{Q}))\geq g(C)-1$ is given in [Gaj22, Theorem 12(1)], whereas $\mathrm{Rank}(\mathrm{Jac}(C)(\mathbb{Q}))\geq g(C)$ is given in [Gaj22, Theorem 12(2)]. The main idea was to find curves with too many rational points that would violate Coleman’s bound [Col85], or Stoll’s improvement [Sto06]. However, giving an exact rank of such curves seems extremely difficult.
In this manuscript, we shift our focus to studying arithmetic statistics of rank of $\mathrm{Jac}(C)(K)$, where $C$ varies over an infinite family of curves. There is a wealth of previous studies which analyse the statistics of rank over infinite family of curves, for example over universal families of elliptic curves [BS15], quadratic twist families of elliptic curves [KMR14, Smi22a, Smi22b, Smi25], cubic twist families of elliptic curves [ABS22, KS24], quartic twist families of elliptic curves over $\mathbb{Q}(i)$ [Sav], and infinitely many elliptic curves of fixed rank between 0 and 4 over any number field [Zyw25a].
1.1. Our results and strategy
The goal of this manuscript is to construct infinitely many non-isomorphic higher genus curves of fixed positive rank.
Notation. Throughout this manuscript, we will not write explicitly the adjective non-isomorphic and the construction of infinitely many curves will always assume they are non-isomorphic. The proof that they are non-isomorphic in each case is straightforward, because we construct the curves by gluing twists of hyperelliptic curves, and for infinitely many twists, we introduce a new prime of bad reduction (using that we twist by square-free elements).
Motivation and difficulties. To the best of our knowledge, the highest exact rank of an elliptic curve over $\mathbb{Q}$ up to this date is 20, found by Elkies and Klagsburn [EK20] (whereas the same people found and elliptic curve of rank at least 29, which is the highest rank elliptic curve over $\mathbb{Q}$, but it is not known if 29 is the exact rank of this curve). Thus, in principle, we can construct an abelian variety over $\mathbb{Q}$ of dimension $g$ and rank $r=20g$, just take $E^{g}$, where $E$ is the known rank 20 curve.
However, there are great mathematical merits to restrict our attention to Jacobians of curves rather than just random abelian varieties because, as we explained before, the curves play a crucial role in investigating many of the most important mathematical questions, and to study and understand their arithmetic, the knowledge of the rank of its Jacobian is one of the most essential information one can obtain. In particular, even though we admit that it is more difficult to compute the rank of a simple Jacobian, as we also mentioned, studying the modular curves lead to so many exciting and groundbreaking results, and they, in principle, have decomposable Jacobians. Hence, it also motivates studying decomposable Jacobians.
In this manuscript, we pay particular focus to studying Jacobians of hyperelliptic curves. Recall that for $g\geq 2$, the dimensions of the moduli space of principally polarised abelian varieties of dimension $g$, the moduli space of curves of genus $g$, and the moduli space of hyperelliptic curves of genus $g$ over $\mathbb{C}$ are respectively given by $\frac{g(g+1)}{2}$, 3g33g-3, and 2g12g-1. For $g\geq 3$, it is clear that the hyperelliptic locus is a thin subset of both moduli spaces of principally polarised abelian varieties and moduli spaces of genus $g$ curves, so the aforementioned strategy of simply taking products of lower dimensional abelian varieties does not construct Jacobians of hyperelliptic curves. Even for $g=2$, it is not clear whether a product of some two elliptic curves is isogenous to the Jacobian of a hyperelliptic curve over a given number field $K$. These considerations highlight one of the subtleties that this manuscript aims to address: we provide explicit gluing procedure for two hyperelliptic curves to produce a new hyperelliptic curve of higher genusdefined over the given number field $K$, thus focusing our problem to thin subset of moduli space of principally polarised abelian varieties.
Another subtlety lies in ensuring that the gluing procedure allows us to incorporate many deep and technical results regarding constructions of hyperelliptic curves whose Jacobians have small fixed rank. Our article demonstrates that all relevant conditions appearing in such previous works can be checked and satisfied even after undertaking the gluing procedure for two hyperelliptic curves over $K$. In particular, we find examples of curves such that effectively, when gluing them, we can consider one of them as a fixed curve (in fact, we use curves which are isomorphic to that one), and then we show that we can apply the known results about the twists of the other one, or we allow twisting both curves with a good control of the rank of both twists concurrently.
Organisation. Observant readers may realise that Section 2 and Section 3 contain similar arguments regarding gluing two hyperelliptic curves to construct a new hyperelliptic curve whose Jacobian has fixed positive rank. The reason why we introduce the genus two case first before the higher genus case in this manuscript is to present our method to potential readers in a more intuitive and accessible manner. Our contributions to this project initially started with genus two hyperelliptic curves. The gluing method is simpler and easier to state for genus 2 hyperelliptic curves, and we hope introducing these techniques first will allow readers to feasibly follow through the generalisations to cases for higher genus hyperelliptic curves.
In Section 2, we consider bielliptic genus two curves $C$ such that $\mathrm{Jac}(C)$ is isogenous to product of two elliptic curves $E_{1}$ and $E_{2}$. We divide the construction into 3 cases. In §2.1, we consider the case over number fields $K$ where both $E_{1}$ and $E_{2}$ do not have any $K$-rational 22-torsion points. In §2.2, we examine elliptic curves over $\mathbb{Q}$ having marked rational $2$-torsion points. In §2.3, we use elliptic curves over $\mathbb{Q}$ having all rational $2$-torsion points. We then construct families of genus two curves $C$ by fixing one curve $E_{2}$ of the prescribed rank $r$, and quadratically twisting $E_{1}$ except in §2.5, which we explain later. This allows us to utilise results from previous literature (such as [Nag29, Fre84, KMR14]) that there are infinitely many such twists of rank 0 or 1 depending on the choice of $E_{1}$. We can hence construct infinitely many bielliptic curves of prescribed rank. For $\boldsymbol{K=\mathbb{Q}}$, we construct the desired families for $\boldsymbol{0\leq r\leq 11}$.
For most of our constructions, it is crucial to fix an elliptic curve of the prescribed rank and to aim for twists of rank 0. Other strategies, for instance, to fix an elliptic curve of rank 0 and to aim for infinitely many twists of rank $r$, seem much more difficult, as in principle, it is much easier to show the existence of infinite families of curves of rank 0 rather than positive rank. We prove that there are not many rational points on these curves, as explained in §2.4. We use positive rank twists in one of the examples in §2.2 and §2.5.
In §2.5, we also construct infinitely many curves over any number field $K$ of genus two, whose rank of the Jacobian over $K$ is each of 0, 1, and 2. These constructions use groundbreaking results in [Smi22b, KP25], which give an existence of infinitely many quadratic twists of a certain elliptic curve over $K$ of rank 0 [Smi22b] or 1 [KP25]. The conditions for elliptic curves for which their theorems can be applied are more complicated than those over $\mathbb{Q}$. Unlike the previous three subsections, the non-zero rank examples we construct crucially depend on rank 1 quadratic twist families. Checking whether the conditions appearing in [KP25] are applicable to our construction of genus two curves is an additional technicality we address in this subsection. The heart of our result for number fields lies in Theorem 2.8, where we propose an intricate construction for such genus two curves. Our work can be considered as a continuation of the recent progress by pioneering constructions of infinitely many elliptic curves of rank exactly equal to $1$ over any number field [KP25, Zyw25b]. More concretely, our work is an application of the results from [KP25] and of the same question for rank 0 [Smi22b].
Generalizing our strategy explored for genus two curves, in Section 3 we consider hyperellitpic curves $C$ of genus $g\geq 3$ such that $\mathrm{Jac}(C)$ is isogenous to product of two Jacobians of hyperelliptic curves $C_{1}$ and $C_{2}$. In §3.1, we construct infinitely many curves over any number field $K$ of any genus at least $3$, whose rank of the Jacobian over $K$ is each of 0, 1, and 2. The overall strategy of the proof remains identical, but we employ different previous results to generate such curves. We still use [Yu19, BM, Smi22b] for the existence of infinitely many quadratic twists of Jacobians of a certain hyperelliptic curve over $K$ of rank 0. But we use the recent breakthrough by Peter Koymans and Adam Morgan [KM] for the existence of infinitely many such quadratic twists of rank $1$. We check in §3.1 that the conditions appearing in [Smi22b] and [KM] are once again applicable to our construction of higher genus curves, as we highlight in Theorem 3.10 and Theorem 3.13. In remaining subsections, we give examples of infinitely many hyperelliptic curves over $\mathbb{Q}$ of genus 3, 4, 5, and 6 with fixed positive rank. To the best of our knowledge, our work is the first attempt to utilize gluing method in the setting of generating infinitely many higher genus hyperelliptic curves with positive fixed rank.
The groundbreaking work by Peter Koymans and Adam Morgan [KM] in fact answers the previously posed challenge problem, which asks for generating infinitely many absolutely simple Jacobians of hyperelliptic curves over $\mathbb{Q}$ with rank $1$. We mention the formulation of the previously posed question in Section 4.
1.2. Explicit families of positive rank
Even though there are several instances of families of curves of fixed positive rank, such as [ABS22, KP25, Zyw25b], not so many give examples of families with explicit parametrisation. There are nevertheless some previous studies which give explicit parametrisation of families of elliptic curves over $\mathbb{Q}$ with rank $1$, such as those proven by Monsky [Mon92] and Dasgupta and Voight [DV18]. Also, very recently, the question of whether there are infinitely many rank 2 elliptic curves over $\mathbb{Q}$ was proved by Zywina [Zyw25c], who, impressively, constructs elliptic curves non-isomorphic over $\overline{\mathbb{Q}}$, and we can say that these results are semi-explicit. While the elliptic curves are given by concrete equations, these equations depend on a certain primality condition that can be viewed as non-explicit, but we still believe that this result deserves to be mentioned in this section. We give explicit examples of curves of genus two and a fixed rank between 0 and 4 in §2.2 and §2.3.
The Magma [BCP97] code for our computations is available at https://github.com/StevanGajovic/Ranks-Weil-Pairings.
2. Genus two curves
We consider bielliptic curves $C$ of genus two; then $\mathrm{Jac}(C)\sim E_{1}\times E_{2}$. In our constructions we fix one curve, say $E_{1}$, and we vary $E_{2}$. Recall that, in general, $E_{1}\times E_{2}$ is not isogenous to a Jacobian of a genus two curve over $K$, but in our cases, the curves $E_{1}$ and $E_{2}$ are “compatible” so it is. More formally, we can glue them so that one gets a bielliptic curve of genus 2, whose Jacobian is isogenous to $E_{1}\times E_{2}$. We provide the constructions for all possible rational two torsion subgroups when considered over $\mathbb{Q}$, i.e., $E_{1}[2](\mathbb{Q})\cong E_{2}[2](\mathbb{Q})\in\{\langle O\rangle,\mathbb{Z}/2\mathbb{Z},\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}\}$. However, our construction in §2.1 works over more general number fields.
2.1. No rational two-torsion
Theorem 2.1.
Let $K$ be a number field and $r\in\mathbb{Z}_{\geq 0}$. Assume that there is an elliptic curve $E\colon y^{2}=x^{3}+a^{2}$, where $a\in\mathcal{O}_{K}\setminus\mathcal{O}_{K}^{3}$ such that $\mathrm{Rank}(E(K))=r$ and $K(E[2])/K$ is an $S_{3}$ extension. There are infinitely many genus two curves $C_{a,m}\colon y^{2}=x^{6}+m^{3}a$ defined over $K$ such that $\mathrm{Rank}(\mathrm{Jac}(C)(K))=r$.
Proof.
We first construct hyperelliptic curves of genus two which are also bielliptic. Given any $D\in K^{\times}$, consider the three curves over $K$:
(1) $\begin{split}C_{D}&\colon y^{2}=x^{6}+D,\\ E_{1,D}&\colon y^{2}=x^{3}+D,\\ E_{2,D}&\colon y^{2}=x^{3}+D^{2}.\end{split}$
Then there exists an isogeny $\phi\colon\mathrm{Jac}(C_{D})\to E_{1,D}\times E_{2,D}$. A direct application of [BD11, Theorem 3.2] shows that $\mathrm{Jac}(C_{D})$ is isogenous to a product of two elliptic curves whose Weierstrass equations are $E^{\prime}_{1,D}\colon y^{2}=x^{3}+D$ and $E^{\prime}_{2,D}\colon y^{2}=x^{4}+Dx$. A change of variables given by $X:=\frac{D}{x}$ and $Y:=\frac{Dy}{x^{2}}$ gives the Weierstrass equation for $E_{2,D}$.
Using these curves, we construct our desired infinite family of hyperelliptic curves as follows. Fix an element $a\in\mathcal{O}_{K}\setminus\mathcal{O}_{K}^{3}$ satisfying the conditions of the statement of the theorem. Let $\mathcal{D}_{a}$ be the collection of elements in $\mathcal{O}_{K}$ defined as
(2) $\mathcal{D}_{a}:=\left\{a\cdot m^{3}\in\mathcal{O}_{K}\colon m\in\mathcal{O}_{K}^{\times}\text{ is square-free}\right\}.$
Then for any $D\in\mathcal{D}_{a}$, the Weierstrass equation for $E_{2,D}$ can be rewritten as $E\colon y^{2}=x^{3}+a^{2}$, which is independent of the choice of $D$. Because $K(E[2])/K$ is an $S_{3}$ extension, $K(E_{1,D}[2])/K$ is also an $S_{3}$ extension. We note that $\{E_{1,D}\}_{D\in D_{a}}$ is a family of quadratic twists of the elliptic curve $E^{\prime}\colon y^{2}=x^{3}+a$. By [KMR14, Theorem A], there exists an infinite subset $\mathcal{D}^{0}_{a}\subset\mathcal{D}_{a}$ such that $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(E_{1,D}/K)=0$, hence $\mathrm{Rank}E_{1,D}(K)=0$, for every $D\in\mathcal{D}^{0}_{a}$. Hence, we have that $\{C_{D}\}_{D\in D^{0}_{a}}$ is an infinite set of hyperelliptic curves such that $\mathrm{Rank}(\mathrm{Jac}(C_{D})(K))=r$. ∎
Corollary 2.2.
Let $0\leq r\leq 11$ be a fixed integer. There are infinitely many genus two curves $C$ defined over $\mathbb{Q}$, such that $\mathrm{Rank}(\mathrm{Jac}(C)(\mathbb{Q}))=r$.
Proof.
For $0\leq r\leq 4$, we apply Theorem 2.1 for the following values of $a=2,3,15,427,13\cdot 19\cdot 23\cdot 43$ for $r=0,1,2,3,4$, respectively. The corresponding rank can easily be checked in Magma [BCP97]. For $5\leq r\leq 11$, we apply the main result from [ER04]. We note that the elliptic curve $y^{2}=x^{3}-432k^{2}$ is $3$-isogenous to the elliptic curve $y^{2}=x^{3}+16k^{2}$ [CP09, Section 1]. We can then set $a=4k$ for the following values of $k$ to produce infinitely many genus two curves of rank $r$ as provided below.
(3) $\begin{split}r=5&:k=3\cdot 7\cdot 11\cdot 13\cdot 163,\\ r=6&:k=3\cdot 73\cdot 103\cdot 439,\\ r=7&:k=3\cdot 13\cdot 19\cdot 41\cdot 139\cdot 271,\\ r=8&:k=2\cdot 3\cdot 5\cdot 7\cdot 11\cdot 13\cdot 17\cdot 29\cdot 41\cdot 47\cdot 59,\\ r=9&:k=2\cdot 5\cdot 37\cdot 41\cdot 53\cdot 73\cdot 1231\cdot 4831,\\ r=10&:k=2\cdot 3\cdot 5\cdot 7\cdot 23\cdot 31\cdot 37\cdot 43\cdot 83\cdot 109\cdot 151\cdot 421,\\ r=11&:k=3\cdot 5\cdot 7\cdot 13\cdot 19\cdot 23\cdot 31\cdot 43\cdot 59\cdot 61\cdot 73\cdot 79\cdot 103\cdot 109\cdot 157\cdot 457.\end{split}$
Remark 2.3.
The condition of $K(E[2])/K$ being an $S_{3}$ extension amounts to $\zeta_{3}\notin K$ (necessary condition) and $\sqrt[3]{a}\notin K$. Therefore, for all number fields $K$ such that $3\nmid[K:\mathbb{Q}]$ and $\zeta_{3}\notin K$, this condition is satisfied. Otherwise, for fields with $\zeta_{3}\notin K$, but with $\sqrt[3]{a}\in K$, one can try to find another elliptic curve $E^{\prime}\colon y^{2}=x^{3}+a^{\prime 2}$ satisfying the conditions of Theorem 2.1.
Remark 2.4.
Suppose $K=\mathbb{Q}$. Using the main results from [Smi22a, Smi22b, Smi25], we can nullify the condition from Theorem 2.1 that $K(E[2])/K$ is an $S_{3}$ extension by allowing $a$ to be a cube. Let $a\in\mathbb{Q}^{\times}$ be a fixed number. As before, let $E:y^{2}=x^{3}+a^{2}$. Then at least $50\%$ of genus two curves in $\{C_{D}\}_{D\in\mathcal{D}_{a}}$ have Jacobians of rank equal to $\mathrm{Rank}(E(\mathbb{Q}))$, and $100\%$ of genus two curves in $\{C_{D}\}_{D\in\mathcal{D}_{a}}$ have Jacobians of rank at most $\mathrm{Rank}(E(\mathbb{Q}))+1$.
2.2. Partial rational two-torsion
Now we consider the family of bielliptic curves $C_{m,d}\colon y^{2}=d^{3}x^{6}+m^{3}$, where $d,m\in\mathbb{Z}$ are square-free. Similarly as before, [BD11, Theorem 3.2] implies that $\mathrm{Jac}(C_{m,d})$ is isogenous to $E_{d}\times E_{m}$, where $E_{d}\colon y^{2}=x^{3}+d^{3}$ and $E_{m}\colon y^{2}=x^{3}+m^{3}$, which is a quadratic twist of $E_{d}$ by $\frac{m}{d}$. Now, we fix $d$ and vary $m$, obtaining the following theorem.
Theorem 2.5.
Let $r\in\mathbb{Z}_{\geq 0}$ and $d\in\mathbb{Z}$ be such that $\mathrm{Rank}(E_{d}(\mathbb{Q}))=r$, where $E_{d}\colon y^{2}=x^{3}+d^{3}$. Then for the following curves of genus two $C_{p,d}\colon y^{2}=d^{3}x^{6}+p^{3}$, where $p$ is a prime number, we have
  • if $p\equiv 5\pmod{12}$, $\mathrm{Rank}(C_{p,d}(\mathbb{Q}))=r$;
  • if $p\equiv 3\pmod{4}$ and $p>3$, $\mathrm{Rank}(C_{p,d}(\mathbb{Q}))=r+1$;
Proof.
The proof follows from the text above and the fact that $\mathrm{Rank}(E_{p}(\mathbb{Q}))=0,1$, when $p\equiv 5\pmod{12}$, $p\equiv 3\pmod{4}$ and $p>3$, respectively, see [Fre84, P. 73, Satz 2, 3]. ∎
We note that this approach works only over $\mathbb{Q}$, but unlike the previous approach, it gives explicit families of infinitely many curves of fixed positive rank. A quick Magma search gives us the following rank information:
  • $d=1$: $E_{1}\colon y^{2}=x^{3}+1$, $\mathrm{Rank}(E_{1}(\mathbb{Q}))=0$;
  • $d=2$: $E_{2}\colon y^{2}=x^{3}+2^{3}$, $\mathrm{Rank}(E_{2}(\mathbb{Q}))=1$;
  • $d=37$: $E_{37}\colon y^{2}=x^{3}+37^{3}$, $\mathrm{Rank}(E_{37}(\mathbb{Q}))=2$;
  • $d=506$: $E_{506}\colon y^{2}=x^{3}+506^{3}$, $\mathrm{Rank}(E_{506}(\mathbb{Q}))=3$.
For example, $y^{2}=506^{3}x^{6}+p^{3}$, where $p\equiv 3\pmod{4}$ and $p>3$, is a family of curves whose Jacobian has rank 4, whereas if $p\equiv 5\pmod{12}$, then this is a family of curves whose Jacobian has rank 3.
2.3. Full rational two-torsion - over $\mathbb{Q}$
Now we consider elliptic curves $E_{1},E_{2}/\mathbb{Q}$ such that $E_{1}[2](\mathbb{Q})\cong E_{2}[2](\mathbb{Q})\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. Since we want to vary one of them, we will specify one of them to be a congruent problem elliptic curve, i.e., let $E_{1}$ belong to the family $E^{p}\colon y^{2}=x^{3}-p^{2}x$, where $p\equiv 3\pmod{8}$; then by [Nag29], we have $\mathrm{Rank}(E^{p}(\mathbb{Q}))=0$. We want to use [BD11, Theorem 3.2], so, we look at the other elliptic curve of the form ($k\neq-1,0,1$)
$E_{2}\colon y^{2}=d(x-kp)(x^{3}-p^{2}x),$
which after replacing $(x,y)$ by $(px,p^{2}y)$ becomes $E_{2,d,k}\colon y^{2}=d(x-k)(x^{3}-x)$. From now on, we can search directly in Magma for curves $E_{2,d,k}$ of any rank we want. Still, if we prefer to use the standard way of expressing elliptic curves, a rational change of variables leads to the following model
$E^{\prime}_{2,d,k}\colon dy^{2}=(x+k^{2}-k)(x+k^{2}-1)(x+k^{2}+k).$
We obtain the following theorem, which immediately follows from [BD11, Theorem 3.2].
Theorem 2.6.
Let $r\in\mathbb{Z}_{\geq 0}$, $k\in\mathbb{Z}\setminus\{-1,0,1\}$, and square-free $d\in\mathbb{Z}$ be such that $\mathrm{Rank}(E^{\prime}_{2,d,k}(\mathbb{Q}))=r$, where $E^{\prime}_{2,d,k}\colon dy^{2}=(x+k^{2}-k)(x+k^{2}-1)(x+k^{2}+k)$. Then, for the following curves of genus two
$C_{d,k,p}\colon dy^{2}=x^{6}+3dkpx^{4}+(3k^{2}p^{2}-1)d^{2}x^{2}+(k^{3}p^{3}-kp)d^{3},$
where $p\equiv 3\pmod{8}$ is a prime number, we have $\mathrm{Rank}(\mathrm{Jac}(C_{d,k,p})(\mathbb{Q}))=r$.
Testing the curves $E^{\prime}_{2,d,k}$ for $d=1$ in Magma, we obtain the following rank information:
  • $k=2$: $\mathrm{Rank}(E^{\prime}_{2,1,2}(\mathbb{Q}))=0$;
  • $k=3$: $\mathrm{Rank}(E^{\prime}_{2,1,3}(\mathbb{Q}))=1$;
  • $k=11$: $\mathrm{Rank}(E^{\prime}_{2,1,11}(\mathbb{Q}))=2$;
  • $k=43$: $\mathrm{Rank}(E^{\prime}_{2,1,43}(\mathbb{Q}))=3$;
  • $k=329$: $\mathrm{Rank}(E^{\prime}_{2,1,329}(\mathbb{Q}))=4$.
For example, one more explicit family of curves of genus two whose Jacobians have rank 4 is the family $C_{1,329,p}$, where $p\equiv 3\pmod{8}$.
2.4. Rational points on these curves
[…]
For example, curves from §2.1 and §2.2 admit a quotient map to an elliptic curve of the form $E\colon y^{2}=x^{3}+s$, $s\in\mathbb{Z}$. By a well-known result [Sil09, Exercise 10.19], when $s\neq-432$ is not a square or a cube in $\mathbb{Z}$, then $E_{\mathrm{tors}}=\langle O\rangle$. This is the case for almost all curves in §2.1, so $C_{a,m}(\mathbb{Q})=\{\infty_{\pm}\}$ (with projective coordinates $(1:\pm 1:0)$).
On the other hand, also by [Sil09, Exercise 10.19], for $E_{p}\colon y^{2}=x^{3}+p^{3}$ with $p\equiv 5\pmod{12}$, we have $E_{p}(\mathbb{Q})=\{O,(-p,0)\}$. Hence, if $d$ in a square (i.e. $d=1$ since we assume $d$ is square-free), then $C_{p,d}(\mathbb{Q})=\{\infty_{\pm}\}$. Otherwise, the set $C_{p,d}(\mathbb{Q})$ is empty. When $p\equiv 3\pmod{4}$, the problem of determining $C_{p,d}(\mathbb{Q})$ could be very difficult.
In §2.3, the structure of the $\mathbb{Q}$-rational points of rank 0 elliptic curve $E^{p}\colon y^{2}=x^{3}-p^{2}x$ is well known: $E^{p}(\mathbb{Q})=E^{p}(\mathbb{Q})[2]\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. Pullback of nontrivial two-torsion points to $C_{d,k,p}$ would give an $x$-coordinate satisfying $x^{2}\in\{-dkp,-d(k-1)p,-d(k+1)p\}$, but then $x\notin\mathbb{Q}$ for almost all $p$. Hence, if $d=1$, then we have $C_{d,k,p}(\mathbb{Q})=\{\infty_{\pm}\}$ almost always. Otherwise, we have $C_{d,k,p}(\mathbb{Q})=\emptyset$.
2.5. Full rational two-torsion - over any number field
Let $K$ be any number field. We now consider elliptic curves $E_{1},E_{2}/K$ such that $E_{1}[2](K)\cong E_{2}[2](K)\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. Unlike the case where $K=\mathbb{Q}$, we will vary both elliptic curves over their families of quadratic twists. Let $a_{1},a_{2},a_{3}\in\mathcal{O}_{K}\setminus\{0\}$. We consider the following elliptic curve, where $m$ is any square-free element of $\mathcal{O}_{K}$.
$E_{1,m}$ $:y^{2}=m(x-a_{1})(x-a_{2})(x-a_{3})$
We then choose $\alpha\in\mathcal{O}_{K}\setminus\{a_{1},a_{2},a_{3}\}$ and a square-free element $d\in\mathcal{O}_{K}$ to define the following elliptic curve by using [BD11, Theorem 3.2]:
$E_{2}:y^{2}=d(x-m\alpha)(x-ma_{1})(x-ma_{2})(x-ma_{3}).$
After replacing $(x,y)$ with $(mx,m^{2}y)$, we obtain $E_{2,d}:y^{2}=d(x-\alpha)(x-a_{1})(x-a_{2})(x-a_{3})$. We can summarize our findings as stated in the following theorem.
Theorem 2.7.
Let $r_{1},r_{2}\in\mathbb{Z}_{\geq 0}$, $a_{1},a_{2},a_{3}\in\mathcal{O}_{K}$, $\alpha\in\mathcal{O}_{K}\setminus\{a_{1},a_{2},a_{3}\}$, and $m,d\in\mathcal{O}_{K}$ be square-free. Let $E_{1,m}:y^{2}=m(x-a_{1})(x-a_{2})(x-a_{3})$ and $E_{2,d}:y^{2}=d(x-\alpha)(x-a_{1})(x-a_{2})(x-a_{3})$ be elliptic curves such that $\mathrm{Rank}(E_{1,m}(K))=r_{1}$ and $\mathrm{Rank}(E_{2,d}(K))=r_{2}$. Then for the following curve of genus two
$C_{m,d}:y^{2}=m\left(\frac{1}{d}x^{2}+\alpha-a_{1}\right)\left(\frac{1}{d}x^{2}+\alpha-a_{2}\right)\left(\frac{1}{d}x^{2}+\alpha-a_{3}\right)$
we have $\mathrm{Rank}(C_{m,d}(K))=r_{1}+r_{2}$.
Combining the above theorem with recent breakthroughs in distribution of algebraic rank of quadratic twist families of elliptic curves over number fields [Smi22b, KP25], we can prove the following result for genus 2 curves over any number field $K$.
Theorem 2.8.
Let $K$ be any number field. Let $0\leq r\leq 2$ be a fixed integer. Then there are infinitely many genus two curves $C$ defined over $K$ such that $\mathrm{Rank}(Jac(C)(K))=r$.
Proof.
We choose nine prime elements $p_{i},q_{i},r_{i}$ of $\mathcal{O}_{K}$ so that the following conditions are satisfied:
  • $p_{i},q_{j},r_{k}$ are totally positive.
  • $p_{i},q_{j},r_{k}\equiv 1\mod 8\mathcal{O}_{K}$.
  • $\mathcal{O}_{K}/(p_{i}),\#\mathcal{O}_{K}/(q_{i}),\#\mathcal{O}_{K}/(r_{i})>5$.
  • For every $1\leq i,j,k\leq 3$, we have $\left(\dfrac{p_{i}}{q_{j}}\right)=\left(\dfrac{q_{j}}{r_{k}}\right)=\left(\dfrac{r_{k}}{p_{i}}\right)=1$.
We will use the following result which is an immediate consequence of [Mit56, P. 20, Theorem]
Lemma 2.9.
Let $I$ be an ideal of $K$ and $\rho\in\mathcal{O}_{K}$ be totally positive such that the ideals $I$ and $(\rho)$ are coprime. Then there are infinitely many totally positive prime elements of $p\in K$ satisfying $p\equiv\rho\pmod{I}$.
Now we apply Lemma 2.9 to the ideal $I=(8)$ and $\rho=1$ to find totally positive prime elements $p_{1},p_{2},p_{3}\in\mathcal{O}_{K}$ such that $\mathcal{O}_{K}/(p_{i})>5$. Then, we apply Lemma 2.9 to the ideal $I=(8p_{1}p_{2}p_{3})$ and $\rho=1$ to find totally positive prime elements $q_{1},q_{2},q_{3}\in\mathcal{O}_{K}$ such that $\mathcal{O}_{K}/(q_{i})>5$. Finally, we apply Lemma 2.9 to the ideal $I=(8p_{1}p_{2}p_{3}q_{1}q_{2}q_{3})$ and $\rho=1$ to find totally positive prime elements $r_{1},r_{2},r_{3}\in\mathcal{O}_{K}$ such that $\mathcal{O}_{K}/(r_{i})>5$. By construction, we have that $\left(\dfrac{p_{i}}{q_{j}}\right)=\left(\dfrac{q_{j}}{r_{k}}\right)=\left(\dfrac{r_{k}}{p_{i}}\right)=1$.
For such a choice of $p_{i},q_{j},r_{k}$, we define $P$, $Q$, $R$ and $a_{1},a_{2},a_{3}$ as follows:
$P:=p_{1}p_{2}p_{3},\hskip 5.0ptQ:=q_{1}q_{2}q_{3},\hskip 5.0ptR:=r_{1}r_{2}r_{3},\hskip 5.0pta_{1}:=P^{2}Q,\hskip 5.0pta_{2}:=Q^{2}R,\hskip 5.0pta_{3}:=R^{2}P.$
Let $m,d$ be square-free elements of $\mathcal{O}_{K}$. We consider the family of genus $2$ curves $C_{m,d}$ defined as
$C_{m,d}:y^{2}=m\left(\frac{1}{d}x^{2}-a_{1}\right)\left(\frac{1}{d}x^{2}-a_{2}\right)\left(\frac{1}{d}x^{2}-a_{3}\right),$
which is isogenous to a product of two elliptic curves provided below:
$E_{1,m}$ $:y^{2}=m(x-a_{1})(x-a_{2})(x-a_{3}),$
$E_{2,d}$ $:y^{2}=d(x-a_{2}a_{3})(x-a_{1}a_{3})(x-a_{1}a_{2}).$
The Weierstrass equation for $E_{2,d}$ is obtained as follows. We set $\alpha=0$ and apply Theorem 2.7 to get $E_{2}:y^{2}=dx(x-a_{1})(x-a_{2})(x-a_{3})$. We then apply the rational change of coordinates which replaces $(x,y)$ with $(-\frac{1}{x}\cdot{a_{1}a_{2}a_{3}},\frac{y}{x^{2}}\cdot a_{1}a_{2}a_{3})$ to get the Weierstrass equation for $E_{2,d}$.
Notice that the four conditions on nine prime elements $p_{i},q_{j},r_{k}$ imply the following (because they all have odd valuation with respect to at least one of the primes $p_{i},q_{j},r_{k}$).
  • None of $(a_{1}-a_{2})(a_{1}-a_{3})$, $(a_{2}-a_{1})(a_{2}-a_{3})$, and $(a_{3}-a_{1})(a_{3}-a_{2})$ are squares in $K$.
  • None of $a_{2}a_{3}(a_{2}-a_{1})(a_{3}-a_{1})$, $a_{1}a_{3}(a_{1}-a_{2})(a_{3}-a_{2})$, and $a_{1}a_{2}(a_{1}-a_{3})(a_{2}-a_{3})$ are squares in $K$.
By [Smi22b, Theorem 2.14, Example 3.1], there exists infinitely many (in fact a positive proportion of) square-free elements $m,d\in\mathcal{O}_{K}$ such that $\mathrm{Rank}(E_{1,m}(K))=\mathrm{Rank}(E_{2,d}(K))=0$. Here, we note that there is a subtlety with generating infinitely many such rank 0 quadratic twist families of elliptic curves over any number field $K$. This is where the “parity-invariant subcase” appearing after [Smi22b, Case 2.13] becomes relevant. (We would like to sincerely thank Adam Morgan for pointing out this subtlety.) Our construction of elliptic curves $E_{1,m}$ and $E_{2,d}$ shows that both curves have at least one place of multiplicative reduction. By the first paragraph of proof of [MS, Proposition 4.3] and [MS, Lemma 4.2], performing an unramified quadratic twist at the local condition over such a multiplicative place changes the parity of 2-Selmer group of $E_{1,m}$ (and likewise for $E_{2,d}$). Hence, we can always find infinitely many $m,d\in\mathcal{O}_{K}$ (in fact a positive density of square-free $m,d\in\mathcal{O}_{K}$) such that the parity of 2-Selmer group of $E_{1,m}$ is even. By [Smi22b, Theorem 2.14, Example 3.1], there are infinitely many choices of $m,d\in\mathcal{O}_{K}$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{m,d})(K))=0$.
We turn our focus to constructing infinitely many curves $C_{m,d}$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{m,d})(K))=1$. We note that this result for $g=2$ was obtained before [KM]. By our choice of nine prime elements $p_{i},q_{j},r_{k}$ the elliptic curve $E_{1,1}$ satisfies the following three conditions:
  • $a_{1}-a_{2}$ has odd valuation with respect to primes $q_{1},q_{2}$, and $q_{3}$, whereas $a_{1}-a_{3}$ and $a_{2}-a_{3}$ are invertible squares modulo $q_{1},q_{2}$, or $q_{3}$.
  • $a_{2}-a_{3}$ has odd valuation with respect to primes $r_{1},r_{2}$, and $r_{3}$, whereas $a_{2}-a_{1}$ and $a_{3}-a_{1}$ are invertible squares modulo $r_{1},r_{2}$, or $r_{3}$.
  • $a_{3}-a_{1}$ has odd valuation with respect to primes $p_{1},p_{2}$, and $p_{3}$, whereas $a_{3}-a_{2}$ and $a_{1}-a_{2}$ are invertible squares $p_{1},p_{2}$, or $p_{3}$.
These three conditions are in fact the “3-generic” conditions appearing in [KP25, Definition 1.2]. By [KP25, Theorem 1.3], there exist infinitely many square-free elements $m$ such that $\mathrm{Rank}(E_{1,m}(K))=1$. We can again use [Smi22b, Theorem 2.14, Example 3.1] to obtain that there are infinitely many square-free elements $d\in\mathcal{O}_{K}$ such that $\mathrm{Rank}(E_{2,d}(K))=0$. For such $m$ and $d$ we have $\mathrm{Rank}(\mathrm{Jac}(C_{m,d})(K))=1$.
Lastly, we construct infinitely many curves $C_{m,d}$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{m,d})(K))=2$. We still choose nine totally positive prime elements $p_{i},q_{j},r_{k}\in\mathcal{O}_{K}$ in the same manner.
We also choose the same $a_{1},a_{2},a_{3}$, but now we change $\alpha$. Consider $\alpha\in\mathcal{O}_{K}$ such that $\alpha$ is a quadratic residue modulo all $p_{i},q_{j},r_{k}$, but such that $\alpha\not\equiv a_{1}\pmod{r_{i}}$, $\alpha\not\equiv a_{2}\pmod{p_{i}}$, and $\alpha\not\equiv a_{3}\pmod{q_{i}}$ for $1\leq i\leq 3$, by the assumption that all residue fields have cardinality greater than 5, we can certainly satisfy the conditions modulo each prime, and then by the Chinese Remainder Theorem obtain $\alpha$.
For such a choice of $p_{i},q_{j},r_{k},$ and $\alpha$, we consider the family of genus 2 curves $C_{m,d}$ defined as
$C_{m,d}:y^{2}=m\left(\frac{1}{d}x^{2}+\alpha-a_{1}\right)\left(\frac{1}{d}x^{2}+\alpha-a_{2}\right)\left(\frac{1}{d}x^{2}+\alpha-a_{2}\right)$
which is isogenous to a product of two elliptic curves provided below:
$E_{1,m}$ $:y^{2}=m(x-a_{1})(x-a_{2})(x-a_{3}),$
$E_{2,d}$ $:y^{2}=d(x+(a_{2}-\alpha)(a_{3}-\alpha))(x+(a_{1}-\alpha)(a_{3}-\alpha))(x+(a_{1}-\alpha)(a_{2}-\alpha)).$
As before, our choices of $p_{i},q_{j},r_{k}$ imply that $E_{1,1}$ is $3$-generic. The choice of $\alpha$ implies that the following three conditions hold:
  • $(\alpha-a_{3})(a_{1}-a_{2})$ has odd valuation with respect to primes $q_{1},q_{2}$, and $q_{3}$, whereas $(\alpha-a_{2})(a_{1}-a_{3})$ and $(\alpha-a_{1})(a_{2}-a_{3})$ are invertible squares modulo $q_{1},q_{2}$, or $q_{3}$.
  • $(\alpha-a_{1})(a_{2}-a_{3})$ has odd valuation with respect to primes $r_{1},r_{2}$, and $r_{3}$, whereas $(\alpha-a_{3})(a_{2}-a_{1})$ and $(\alpha-a_{2})(a_{3}-a_{1})$ are invertible squares modulo $r_{1},r_{2}$, or $r_{3}$.
  • $(\alpha-a_{2})(a_{3}-a_{1})$ has odd valuation with respect to primes $p_{1},p_{2}$, and $p_{3}$, whereas $(\alpha-a_{1})(a_{3}-a_{2})$ and $(\alpha-a_{3})(a_{1}-a_{2})$ are invertible squares modulo $p_{1},p_{2}$, or $p_{3}$.
These three conditions show that $E_{2,1}$ is 3-generic. By [KP25, Theorem 1.3], there exist infinitely many square-free elements $m$ and $d$ such that $\mathrm{Rank}(E_{1,m}(K))=1$ and $\mathrm{Rank}(E_{1,d}(K))=1$. For such $m$ and $d$ we have $\mathrm{Rank}(\mathrm{Jac}(C_{m,d})(K))=2$. ∎
3. Higher genus curves
We generalise the method of gluing two elliptic curves explored in Section 2 to gluing two hyperelliptic curves. This allows us to construct infinitely many hyperelliptic curves of any genus over any number field whose Jacobians have fixed positive rank between 0 and $2$. We also explore examples of genus 3, 4, 5, and 6 hyperelliptic curves over $\mathbb{Q}$ with fixed small positive rank.
3.1. Higher genus - over any number field
Let $K$ be a number field, $f=x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0}\in\mathcal{O}_{K}[x]$ a monic polynomial of degree $n$, and $d\in\mathcal{O}_{K}$. Define
$f_{d}(x)=x^{n}+da_{n-1}x^{n-1}+\cdots+d^{n-1}a_{1}x+d^{n}a_{0},$
i.e., the polynomial such that $f_{d}(dx)=d^{n}f(x)$ Consider the following three hyperelliptic curves
$C$ $\colon dy^{2}=f_{d}(x^{2});$
$C_{1}$ $\colon dy^{2}=xf_{d}(x);$
$C_{2}$ $\colon dy^{2}=f_{d}(x).$
The curve $C$ admits two degree 2 morphisms
$\psi_{1}\colon C\longrightarrow C_{1},$ $\;\;\psi_{1}(x,y)=(x^{2},xy);$
$\psi_{2}\colon C\longrightarrow C_{2},$ $\;\;\psi_{2}(x,y)=(x^{2},y).$
Lemma 3.1.
The map $\Psi=({\psi_{1}}_{*},{\psi_{2}}_{*})\colon\mathrm{Jac}(C)\longrightarrow\mathrm{Jac}(C_{1})\times\mathrm{Jac}(C_{2})$ is an isogeny defined over $K$.
Proof.
The map $\Psi$ is defined over $K$ as $\psi_{1},\psi_{2}$ are defined over $K$. We denote by $C^{(g(C))}$, $C_{1}^{(g(C_{1}))}$, and $C_{2}^{(g(C_{2}))}$ the symmetric powers of $C$, $C_{1}$, and $C_{2}$ by respective genus g(C)=n1g(C)=n-1, $g(C_{1})=\lfloor\frac{n}{2}\rfloor$ and $g(C_{2})=\lfloor\frac{n-1}{2}\rfloor$. Using [Mil86b, Theorem 5.1, Corollary 7.5], we have the following commutative diagram.
$C^{g}$$C_{1}^{g(C_{1})}\times C_{2}^{g(C_{2})}$$C^{(g(C))}$$C_{1}^{(g(C_{1}))}\times C_{2}^{(g(C_{2}))}$$\mathrm{Jac}(C)$$\mathrm{Jac}(C_{1})\times\mathrm{Jac}(C_{2})$$\scriptstyle{\psi_{1}\times\psi_{2}}$$\scriptstyle{\Psi}$
The surjectivity of $\Psi$ follows from the surjectivity of $C^{g}\rightarrow C_{1}^{g(C_{1})}\times C_{2}^{g(C_{2})}$ and all vertical maps in the commutative diagram. Because $\Psi$ is a homomorphism of abelian varieties and since $\lfloor\frac{n}{2}\rfloor+\lfloor\frac{n-1}{2}\rfloor=n-1$, we have $\dim\mathrm{Jac}(C)=\dim(\mathrm{Jac}(C_{1})\times\mathrm{Jac}(C_{2}))$, so, we can use [Mil86a, Proposition 8.1] to show that $\Psi$ is an isogeny. ∎
Remark 3.2.
Thanks to Maarten Derickx’s suggestion, we can note that the pull-backs $\psi_{1}^{*}(\mathrm{Jac}(C_{1}))\subseteq\mathrm{Jac}(C)$ and $\psi_{2}^{*}(\mathrm{Jac}(C_{2}))\subseteq\mathrm{Jac}(C)$ have 0-dimensional intersection inside $\mathrm{Jac}(C)$. This can be checked by pulling-back the holomorphic differentials $\frac{dx}{y},\ldots\frac{x^{\lfloor\frac{n}{2}\rfloor-1}dx}{y}$ from $\mathrm{Jac}(C_{1})$ and $\frac{dx}{y},\ldots\frac{x^{\lfloor\frac{n-1}{2}\rfloor-1}dx}{y}$ from $\mathrm{Jac}(C_{2})$, which give the full basis of the holomorphic differentials $\frac{dx}{y},\ldots\frac{x^{n-2}dx}{y}$ on $\mathrm{Jac}(C)$.
Note that, depending on the parity of $n$, after a linear change of variables ($x\mapsto dx$), we can simplify the equations for $C_{1}$ and $C_{2}$
$C_{1}\cong y^{2}=xf(x),\;C_{2}\cong dy^{2}=f(x)$ $\text{ if }2\mid n,$
$C_{1}\cong dy^{2}=xf(x),\;C_{2}\cong y^{2}=f(x)$ $\text{ if }2\nmid n.$
Note that in both cases, one of the curves is isomorphic to a fixed one, given by an odd degree model, whereas an even degree model becomes a twist by $d$. Hence, this gives a clear strategy to construct infinitely many curves of genus $g$ of a fixed rank $r$; we just need to find a specific one satisfying some mild conditions.
Before we proceed, we recall a few definitions and results from previous literature.
Definition 3.3 (Definition 1.2 and 2.2 of [KM]).
Let $C:y^{2}=(x-a_{1})\cdots(x-a_{2g+1})$ be an odd degree hyperelliptic curve over a number field $K$. Given a place $\omega$ of $K$ and a pair of distinct indices $1\leq i,j\leq 2g+1$, we say that $\omega$ is multiplicative of type $\{i,j\}$ if for every pair of distinct indices $1\leq k,\ell\leq 2g+1$, we have
(4) $\mathrm{val}_{\omega}(a_{k}-a_{\ell})=\begin{cases}1&\text{ if }\{k,\ell\}=\{i,j\},\\ 0&\text{ otherwise}.\end{cases}$
Given a positive number $B>0$, we say that $C$ is $B$-generic if there exist $4g^{2}+4g-1$ distinct finite places
$\begin{cases}\omega_{1},\omega_{2},&\cdots,\omega_{4g^{2}+2g-1},\\ \omega^{\prime}_{1},\omega^{\prime}_{2},&\cdots,\omega^{\prime}_{2g}.\end{cases}$
which satisfy the following four conditions.
  • For every $1\leq i,j\leq 2g$, the place $\omega_{(i-1)2g+j}$ is multiplicative of type $\{j,2g+1\}$.
  • For every $1\leq i\leq 2g-1$, the place $\omega_{4g^{2}+i}$ is multiplicative of type $\{i,i+1\}$.
  • For every $1\leq i\leq 4g^{2}+2g-1$, we have $\mathcal{O}_{K}/(\omega_{i})\geq B$.
  • For every $1\leq i\leq 2g$, the place $\omega^{\prime}_{i}$ is of type $\{i,2g+1\}$.
Definition 3.4.
Using the notation of Definition 3.3, we say that $C$ is BB-über-generic if the following conditions are satisfied: There are $g^{2}(2g+1)$ distinct places of $K$ such that for each pair $1\leq k<\ell\leq 2g+1$ we have for exactly $2g+1$ places $\omega$ that
(5) $\mathrm{val}_{\omega}(a_{k}-a_{\ell})=\begin{cases}1&\text{ if }\{k,\ell\}=\{i,j\},\\ 0&\text{ otherwise},\end{cases}$
and for all places $\omega$, and all $a_{i}$, we have $\mathrm{val}_{w}(a_{i})=0$ and $\mathcal{O}_{K}/(\omega)\geq B$.
We focus on understanding the Mordell-Weil rank of the family of quadratic twists of hyperelliptic curves that are BB-über-generic. Before doing so, we recall the following condition on $C$ from [BM, Condition 5.1] and [MS, Proposition 4.3]. Once again, we would like to thank Adam Morgan for suggesting the importance of the third condition provided below.
Condition 3.5 (Condition 5.1 from [BM] and Proposition 4.3 from [MS]).
Let $C$ be a hyperelliptic curve over $K$ with $K(\mathrm{Jac}(C)[2])=K$. Then $C$ satisfies the following three conditions.
  • The invariant subspaces of $\mathrm{Jac}(C)[2]$ with respect to $\mathrm{Gal}(K(\mathrm{Jac}(C)[4])/K)$ are 0 or $\mathrm{Jac}(C)[2]$.
  • The set of elements of $\mathrm{End}$ which commute with every element of $\mathrm{Gal}(K(\mathrm{Jac}(C)[4])/K)$ via multiplication is isomorphic to $\mathbb{F}_{2}$.
  • There exists a place $\omega$ of $K$ such that $C$ has a single nodal singularity at $\omega$. (In particular, the latter condition is satisfied if there exists a place $\omega$ and a tuple $(i,j)$ such that $\mathrm{val}_{\omega}(a_{i}-a_{j})=1$ and $\mathrm{val}_{\omega}(a_{k}-a_{l})=0$ for any other $k,l$ such that $\{k,l\}\neq\{i,j\}$.)
Remark 3.6.
We note that if $C$ is $B$-generic, then there exists a square-free element $d\in\mathcal{O}_{K}^{\times}$ such that $C_{d}$ satisfies Condition 3.5. To see this, we note that [BM, Proposition 8.10] shows that any hyperelliptic curve $C:y^{2}=(x-a_{1})\cdots(x-a_{2g+1})$ satisfying the following two conditions
  • There exists a set of places $\{\omega_{1},\omega_{2},\cdots,\omega_{2g}\}$ such that for each $1\leq i\leq 2g$, $\mathrm{val}_{\omega_{i}}(a_{i}-a_{2g+1})=1$ whereas $\mathrm{val}_{\omega_{i}}(a_{k}-a_{l})=0$ for any other $k,l$ such that $\{k,l\}\neq\{i,2g+1\}$.
  • There exists an element $\zeta\in\mathrm{Sel}_{2}(\mathrm{Jac}(C))\setminus\mathrm{Jac}(C)[2]$ such that the restriction of $\zeta$ over $H^{1}(K_{\omega_{i}},\mathrm{Jac}(C)[2])$ is unramified for all $1\leq i\leq 2g$.
satisfies the first two requirements in Condition 3.5. By [KM, Theorem 3.8], there exists a square-free element $d\in\mathcal{O}_{K}^{\times}$ such that the quadratic twist $C_{d}$ of a $B$-generic hyperelliptic curve $C$ satisfies the two bullet points mentioned above. By the condition on $a_{i}$’s, one can easily check that $C$ has a single nodal singularity at the place $\omega_{1}$, which satisfies the last condition from Condition 3.5.
Remark 3.7.
Let $C$ be a hyperelliptic curve over $\mathbb{Q}$. One can check the first two conditions from Condition 3.5 by computing the Weil pairing $e_{2}:\mathrm{Jac}(C)[2]\times\mathrm{Jac}(C)[2]\to\mu_{2}$. Suppose $\langle P_{1},P_{2},\cdots,P_{2g}\rangle$ is a basis for the set $\mathrm{Jac}(C)[2]$. If for every pair $1\leq i<j\leq 2g$ we have $e_{2}(P_{i},P_{j})=-1$, then the first two conditions hold by [BM, Proposition 2.15].
For the special case when $C$ is given by the Weierstrass model $C:y^{2}=(x-b_{1})\cdots(x-b_{4k+3})$ for some positive integer $k$, we can combine [Yu16, Corollary 7.6] and [Smi22b, Theorem 2.14] to ensure that at least $50\%$ of Jacobians of quadratic twists of $C$ have rank 0. In the case when $C$ is not of the form above, we can replace the third condition in Condition 3.5 with the requirement that there exists a single square-free element $d\in\mathbb{Q}^{\times}$ such that $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(\mathrm{Jac}(C_{d})/\mathbb{Q})\equiv 0\mod 2$. This condition ensures that there is a positive density of Jacobians of rank 0, as shown in the discussion following after [Smi22b, Case 2.13] and the main theorem of [BM].
Theorem 3.8.
Let $C:y^{2}=(x-a_{1})\cdots(x-a_{2g+1})$ be a hyperelliptic curve over a number field $K$ that satisfies Condition 3.5. Then there are infinitely many square-free elements $d\in\mathcal{O}_{K}^{\times}$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{d})(K))=0$.
Proof.
The proof of [KM, Corollary 6.15] and the first paragraph of the proof of [MS, Proposition 4.3] implies that all the conditions in [Smi22b, Case 2.13] are satisfied, and that a positive density of square-free elements $d\in\mathcal{O}_{K}$ satisfies $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(\mathrm{Jac}(C_{d})/K)\equiv 0\mod 2$. Hence, we can apply [Smi22b, Theorem 2.14] to show that there are infinitely many quadratic twists of $C$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{d})/K)=0$. ∎
Corollary 3.9.
For any $g\geq 2$, there are infinitely many hyperelliptic curves $C$ of genus $g$ over a number field $K$ such that $\mathrm{Rank}(\mathrm{Jac}(C)(K))=0$.
Proof.
We apply Theorem 3.8 to any $B$-generic hyperelliptic curve $C$ of genus $g$, which can easily be constructed using the Chinese Remainder Theorem. ∎
Theorem 3.10.
Let $g\geq 3$ be a positive integer and denote $k=\lfloor\frac{g+1}{2}\rfloor$. Assume there are $b_{1},\ldots,b_{2k+1}\in\mathcal{O}_{K}$ so that the hyperelliptic curve
$X\colon y^{2}=(x-b_{1})\cdots(x-b_{2k+1})$
has $\mathrm{Rank}(\mathrm{Jac}(X)(K))=r$. Suppose for each specified value of $g$ as mentioned below, the curve $Y_{1}$ satisfies Condition 3.5:
  • $g=2k$, and $Y_{1}:y^{2}=x(x-b_{1})\cdots(x-b_{2k+1})$.
  • g=2k1g=2k-1, and $Y_{1}:y^{2}=(x-(b_{1}-b_{2k+1}))\cdots(x-(b_{2k}-b_{2k+1}))$.
Then, there are infinitely many hyperelliptic curves of genus $g$ whose Jacobian over $K$ has rank $r$.
Proof.
Suppose $g=2k$, Let $Y_{d}\colon dy^{2}=x(x-b_{1})\cdots(x-b_{2k+1})$ be a quadratic twist of $Y_{1}$ by $d$. By Theorem 3.8, there are infinitely many such quadratic twists $Y_{d}$ such that $\mathrm{Rank}(\mathrm{Jac}(Y_{d})(K))=0$. Hence, there are infinitely many genus $g$ curves
$C_{d}\colon dy^{2}=(x^{2}-db_{1})\cdots(x^{2}-db_{2k+1})$
of rank $r$ as $\mathrm{Jac}(C_{d})$ is isogenous to $\mathrm{Jac}(X)\times\mathrm{Jac}(Y_{d})$ by Lemma 3.1.
Similarly, if g=2k1g=2k-1, then,
$X\cong X^{\prime}\colon y^{2}=x(x-(b_{1}-b_{2k+1}))\cdots(x-(b_{2k}-b_{2k+1})).$
By Theorem 3.8, there are infinitely many twists
$Y_{d}\colon dy^{2}=(x-(b_{1}-b_{2k+1}))\cdots(x-(b_{2k}-b_{2k+1}))$
of rank 0. By Lemma 3.1, there are infinitely many curves
$C_{d}\colon dy^{2}=(x^{2}-d(b_{1}-b_{2k+1}))\cdots(x^{2}-d(b_{2k}-b_{2k+1}))$
of genus $g$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{d})(K))=r$. ∎
Remark 3.11.
We note that the affine models for curves $Y_{1}$ in both cases $g=2k$ and g=2k1g=2k-1 are defined in terms of even degree polynomials in $x$. For computational purposes, we rewrite the models for $Y_{1}$ in terms of odd degree polynomials in $x$. For the case when $g=2k$, let $\mathcal{B}:=\prod_{i=1}^{2k+1}b_{i}$. Then the equation defining $Y_{1}$ can be rewritten as
$Y_{1}\cong Y_{1}^{\prime}:y^{2}=\left(x+\frac{\mathcal{B}}{b_{1}}\right)\cdot\left(x+\frac{\mathcal{B}}{b_{2}}\right)\cdots\left(x+\frac{\mathcal{B}}{b_{2k+1}}\right).$
If g=2k1g=2k-1, let $\mathcal{B}:=\prod_{i=2}^{2k}(b_{i}-b_{1})$. Then we can rewrite the equation defining $Y_{1}$ as
$Y_{1}\cong Y_{1}^{\prime}:y^{2}=\left(x+\frac{\mathcal{B}^{\prime}}{b_{2}-b_{1}}\right)\cdot\left(x+\frac{\mathcal{B}^{\prime}}{b_{3}-b_{1}}\right)\cdots\left(x+\frac{\mathcal{B}^{\prime}}{b_{2k}-b_{1}}\right).$
Then, in our applications, see §3.2, we check Condition 3.5 directly for these odd degree models using Magma.
We can construct such examples in a different way, i.e., coming of the hyperelliptic curves, whose Jacobians have very few $K$-rational 2-torsion points, using the result of Yu, [Yu19, Theorem 1.2].
Theorem 3.12.
Let $g\geq 3,r\geq 0$ be integers and $K$ a number field.
  1. (1)
    Assume that $g=2k$ for some $k\in\mathbb{N}$ and that there exists a hyperelliptic curve $X:y^{2}=f(x)$ such that $\mathrm{Rank}(\mathrm{Jac}(X)(K))=r$, where $f\in\mathcal{O}_{K}[x]$, $\deg(f)=2k+1$, is a monic polynomial whose splitting field has Galois group isomorphic to $S_{2k+1}$. Suppose in addition that there exists a square-free element $m\in\mathcal{O}_{K}$ such that the curve $Y_{m}:my^{2}=xf(x)$ satisfies $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(\mathrm{Jac}(Y_{m})/K)\equiv 0\mod 2$. Then, there are infinitely many genus $g$ curves $C_{d}:dy^{2}=f_{d}(x^{2})$ of rank $r$, where $d\in\mathcal{O}_{K}$.
  2. (2)
    Assume that g=2k1g=2k-1 for some $k\in\mathbb{N}$ and that there exists a hyperelliptic curve $X:y^{2}=x(x-1)f(x)$ such that $\mathrm{Rank}(\mathrm{Jac}(X)(K))=r$, where $f\in\mathcal{O}_{K}[x]$, $\deg(f)=2k-1$, is a monic polynomial whose splitting field has Galois group isomorphic to $S_{2k-1}$. Suppose in addition that there exists a square-free element $m\in\mathcal{O}_{K}$ such that the curve $Y_{m}:my^{2}=(x-1)f(x)$ satisfies $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(\mathrm{Jac}(Y_{m})/K)\equiv 0\mod 2$. Then, there are infinitely many many genus $g$ curves $C_{d}:dy^{2}=(x^{2}-d^{2})f_{d}(x^{2})$ of rank $r$, where $d\in\mathcal{O}_{K}$.
Proof.
  1. (1)
    We note that the existence of a square-free element $m\in\mathcal{O}_{K}$ such that $Y_{m}:my^{2}=xf(x)$ satisfies $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(\mathrm{Jac}(Y_{m})/K)\equiv 0\mod 2$ implies that the disparity $\delta(\mathrm{Jac}(Y_{1})/K)\neq-\frac{1}{2}$ appearing in the statement of [Yu19, Theorem 1.2]. The assumptions assure that we can apply [Yu19, Theorem 1.2] to ensure that there are infinitely many quadratic twists $Y_{d}\colon dy^{2}=xf(x)$ of $Y\colon y^{2}=xf(x)$ such that $\mathrm{Rank}(\mathrm{Jac}(Y_{d})(K))=0$. Then, the statement follows from Lemma 3.1 which gives that $\mathrm{Jac}(C_{d})$ is isogenous to $\mathrm{Jac}(X)\times\mathrm{Jac}(Y_{d})$ over $K$, and the argument that there are infinitely many curves that are non-isomorphic among them is as before.
  2. (2)
    The same as (1), and now we apply [Yu19, Theorem 1.2] to conclude that there are infinitely many quadratic twists $Y_{d}\colon dy^{2}=(x-1)f(x)$ of $Y\colon y^{2}=(x-1)f(x)$ such that $\mathrm{Rank}(\mathrm{Jac}(Y_{d})(K))=0$.
Theorem 3.13.
Let $k$ be a positive integer. There is $B>0$ such that the following holds. Let $X:y^{2}=(x-a_{1})\cdots(x-a_{2k+1})$ be a hyperelliptic curve over a number field $K$ that is BB-über-generic such that $\mathrm{Rank}(\mathrm{Jac}(C)(K))=r$. Then there are infinitely many hyperelliptic curves $X/K$ of genus $g$, where $g\in\{2k-1,2k\}$ satisfying $\mathrm{Rank}(\mathrm{Jac}(C)(X))=r+1$.
Proof.
The easier case is when g=2k1g=2k-1. We may assume that $a_{2k+1}=0$ (otherwise just replace $x-a_{2k+1}$ by $x$). We need to glue the curve $C$ with $Y:y^{2}=(x-a_{1})\cdots(x-a_{2k})$. We choose $B>0$ from [KM, Theorem 1.3] for genus k1k-1. Our assumption of being BB-über-generic clearly implies that $Y$ is $B$-generic. Hence, by [KM, Theorem 1.3], there are infinitely many twists $Y_{d}:dy^{2}=(x-a_{1})\cdots(x-a_{2k})$ of $Y$ such that $\mathrm{Rank}(\mathrm{Jac}(Y_{d})(K))=1$. Then, for $C_{d}:dy^{2}=(x^{2}-da_{1})\cdots(x^{2}-da_{2k})$, by Lemma 3.1, we have that $\mathrm{Jac}(C_{d})$ is isogenous to $\mathrm{Jac}(X)\times\mathrm{Jac}(Y_{d})$ over $K$. Hence, $\mathrm{Rank}(\mathrm{Jac}(C_{d})(K))=r+1$, and the argument that there are infinitely many non-isomorphic curves among them is as before.
Now, we consider the case $g=2k$. We may assume now $a_{1}\cdots a_{2k+1}\neq 0$. We need to glue the curve $C$ with $Y:y^{2}=x(x-a_{1})\cdots(x-a_{2k+1})$. We choose $B>0$ from [KM, Theorem 1.3] for genus $k$. We need an odd model of $Y$, so the change of variables $x\mapsto\frac{1}{x}$, induces that $Y$ is isomorphic over $K$ to
$Z:y^{2}=\left(x+\frac{A}{a_{1}}\right)\cdots\left(x+\frac{A}{a_{2k+1}}\right),$
where $A=a_{1}\cdots a_{2k+1}$. Now, the assumption that $X$ is BB-über-generic gives that for any $i,j$ and valuation $w$
$\mathrm{val}_{w}\left(\frac{A}{a_{i}}-\frac{A}{a_{j}}\right)=\mathrm{val}_{w}\left(\frac{A}{a_{i}a_{j}}(a_{j}-a_{i})\right)=\mathrm{val}_{w}(a_{j}-a_{i}),$
so $Z$ is $B$-generic. Therefore, we can apply [KM, Theorem 1.3] to conclude that there are infinitely many twists $Z_{d}$ of $Z$ such that $\mathrm{Rank}(\mathrm{Jac}(Z_{d})(K))=1=\mathrm{Rank}(\mathrm{Jac}(Y_{d})(K))$ for the corresponding twist $Y_{d}$ of $Y$. Now, we conclude as in the previous case. ∎
Remark 3.14.
Using the same idea of the proof that $X$ being BB-über-generic implies $Y_{d}$ being $B$-generic, we can use Remark 3.6 and [Smi22b, Theorem 2.14] to prove that if $X$ is BB-über-generic and $\mathrm{Rank}(\mathrm{Jac}(X)(K))=r$, then there are infinitely many hyperelliptic curves of genus $g$ whose Jacobian over $K$ has rank $r$.
Theorem 3.15.
Let $g\geq 3$ be an integer and $K$ a number field. There are infinitely many hyperelliptic curves $C$ such that $\mathrm{Rank}(\mathrm{Jac}(C)(K))=r$ for $r=1$ and $2$.
Proof.
By [KM], we can always find a quadratic twist of a BB-über-generic curve $X$ of genus $\lfloor\frac{g+1}{2}\rfloor$ whose Jacobian over $K$ has rank $1$. We then apply Theorem 3.10 (++ Remark 3.6) and Theorem 3.13 to conclude. ∎
3.2. Examples over $\mathbb{Q}$
Now we prove the existence of infinitely many curves over $\mathbb{Q}$ of genus $g$ whose Jacobian over $\mathbb{Q}$ as rank $r$ for $g\in\{3,4\}$ and $1\leq r\leq 4$ and $g\in\{5,6\}$ and $1\leq r\leq 3$. Even though the existence of the rank 1 and 2 follows from Theorem 3.15, we still decided to include these examples.
We will always use Theorem 3.10. We will find a hyperelliptic curve of the shape
$X\colon y^{2}=(x-b_{1})\cdots(x-b_{2k+1})$
such that $\mathrm{Rank}(\mathrm{Jac}(X)(\mathbb{Q}))=r$ and check that the conditions for the curve $Y_{1}$ from Theorem 3.10 using Remark 3.7. We note that only in the case for genus three curves, we do not need to do any checks because in this case $Y_{1}$ is an elliptic curve and then the result about twisting comes automatically from [Smi25].
3.2.1. Genus three and four
We present a table of genus two curves $X\colon y^{2}=(x-b_{1})\cdots(x-b_{5})$, which, by Remark 3.7 and Theorem 3.10, prove the existence of infinitely many hyperelliptic curves $C/\mathbb{Q}$ of genus three and four such that $\mathrm{Rank}(\mathrm{Jac}(C)(\mathbb{Q}))=r$, for the following $r$ from the table.
$r$ $b_{1}$ $b_{2}$ $b_{3}$ $b_{4}$ $b_{5}$
1 1 2 3 4 10
2 1 2 3 4 37
3 -49 -4 1 43 51
4 -49 1 2 28 51
3.2.2. Genus five and six
We present a table of genus three curves $X\colon y^{2}=(x-b_{1})\cdots(x-b_{7})$, which, by Remark 3.7 and Theorem 3.10, prove the existence of infinitely many hyperelliptic curves $C/\mathbb{Q}$ of genus five and six such that $\mathrm{Rank}(\mathrm{Jac}(C)(\mathbb{Q}))=r$, for the following $r$ from the table.
$r$ $b_{1}$ $b_{2}$ $b_{3}$ $b_{4}$ $b_{5}$ $b_{6}$ $b_{7}$
1 1 2 3 4 5 6 7
2 1 2 3 4 6 7 11
3 1 2 3 4 5 14 21
Remark 3.16.
Even though it sounds appealing to use Theorem 3.13 to construct examples, we note that the $B$-generic and BB-über-generic conditions are less practical to use. Namely, a careful inspection on the condition on $B$ from [KM] shows that $B>9600$ for genus two curves, so we will need to use several large primes to construct such $B$-generic or BB-übergeneric curves. Moreover, solving the Chinese Reminder Theorem system will result in large coefficients of the curve $C$. Hence, we have not tried to use $B$-generic or BB-übergeneric curves to get concrete examples.
Also, Theorem 3.12 is more tricky for applications because it is much easier to do explicit descent for elliptic curves defined by a polynomial which splits completely, rather than of those whose Galois group is as maximal as possible. Hence, we have not tried to use this theorem to construct examples either.
4. An answer to our question
If the Jacobian of a curve is decomposable as a product of Jacobians of curves, then we reduce the problem of computing the rank to smaller genus curves, which is, in principle, way easier. The problem becomes much harder if we restrict the question for curves $C$ of any genus such that $\mathrm{Jac}(C)$ is absolutely simple. The groundbreaking work of Peter Koymans and Adam Morgan [KM] confirms that there are infinitely many absolutely simple Jacobians of hyperelliptic curves such that the rank of $\mathrm{Jac}(C)(\mathbb{Q})$ is equal to $1$. This result resolves the following question, which was posed as a challenge in the previous version of this article as stated in June 2025.
Question 4.1.
Is there an infinite family of curves $C/\mathbb{Q}$ of genus two or three with absolutely simple Jacobian such that the rank of $\mathrm{Jac}(C)(\mathbb{Q})$ is equal to $1$ (or some other fixed positive integer)?
We keep the question as stated in its original form, as [KM, Theorem 1.1] specifically cites the problem mentioned beforehand. We sincerely congratulate and thank them for solving the question stated above.
Acknowledgements
We would like to thank the Max Planck Institute for Mathematics for its generous support. We would like to also thank Lodha Mathematical Sciences Institute for its generous support and hospitality during the Thematic Programme on Arithmetic Statistics where substantial results of this paper were achieved. We would especially like to thank Jennifer Balakrishnan, Maarten Derickx, Dušan Dragutinović, Peter Koymans, Adam Morgan, Carlo Pagano, and Lazar Radičević for numerous conversations and their enlightening comments about this work. We would like also to thank Jordan Ellenberg, Matija Kazalicki, Jef Laga, Filip Najman, Oana Padurariu, Alex Smith, and Wadim Zudilin for helpful discussions.